### v2

Abstract
Speculative decoding in large language models (LLMs) accelerates token generation by speculatively predicting multiple tokens cheaply and verifying them in parallel, and has been widely deployed. In this paper, we provide the first study demonstrating the privacy risks of speculative decoding. We observe that input-dependent patterns of correct and incorrect predictions can be leaked out to an adversary monitoring token generation times and packet sizes, leading to privacy breaches. By observing the pattern of correctly and incorrectly speculated tokens, we show that a malicious adversary can fingerprint queries and learn private user inputs with more than 90% accuracy across three different speculative decoding techniques – REST (almost 100% accuracy), LADE (up to 92% accuracy), and BiLD (up to 95% accuracy). We show that an adversary can also leak out confidential intellectual property used to design these techniques, such as data from data-stores used for prediction (in REST) at a rate of more than 25 tokens per second, or even hyper-parameters used for prediction (in LADE). We also discuss mitigation strategies, such as aggregating tokens across multiple iterations and padding packets with additional bytes, to avoid such privacy or confidentiality breaches.
1 Introduction
Large Language Models (LLMs) have transformed natural language processing (NLP), enabling machines to generate and understand human language at an unprecedented scale Vaswani et al. (2017); Devlin et al. (2019); Brown et al. (2020); Zhang et al. (2022); Le Scao et al. (2022); Touvron et al. (2023). LLMs typically generate text using auto-regressive decoding Touvron et al. (2023); OpenAI (2024a), where the generation happens serially and each token depends on all the previous ones. Unfortunately, this serial process causes a significant bottleneck in the LLM response latency Miao et al. (2024); Fu et al. (2024) and under-utilizes the available hardware-level parallelism, limiting token generation throughput and latency.
Speculative decoding Leviathan et al. (2023); Chen et al. (2023); Miao et al. (2024); Spector & Re (2023) addresses this problem without impacting model accuracy. It uses smaller models or heuristics such as retrieval or self-drafting to inexpensively generate tokens speculatively, which the larger target model verifies in parallel in a single iteration. By tuning the heuristics to maintain a high rate of correct speculations, such techniques provide 2x to 5x speedups in inference latency and throughput Xia et al. (2024).
However, the adoption of speculative techniques is not without risks. Speculative execution in CPUs  Burton (1985); Hennessy & Patterson (2012), which inspired speculative decoding, has led to considerable security vulnerabilities in processors, such as Spectre Kocher et al. (2019), Meltdown Lipp et al. (2018), and their numerous variants. These side-channel vulnerabilities exploit the side-effects of processor mis-speculations, such as timing variations introduced by mis-speculations, to leak private and confidential data that become accessible to a processor during mis-speculations. This raises the question, does speculative decoding in LLMs also introduce new risks to privacy?
In this paper, we provide the first study of the privacy and confidentiality risks of speculative decoding in LLMs, including leakage of private user inputs or confidential data.
Problem. We observe two key issues in speculative decoding implementations: (1) the pattern of correct and incorrect speculations of output tokens is dependent on the input, and (2) input-dependent speculation patterns can be inferred from the variations in packet sizes: upon correct speculation, more tokens generated per iteration leads to a larger packet and so observing the packet size leaks the degree of correct speculation. Consequently, an adversary capable of measuring the number of tokens generated per iteration or packet sizes can gain access to input-dependent speculation-patterns based on variation in packet sizes and leak out private input attributes, or even entire inputs and outputs.
Q. How to fine-tune a large language model?   A.  F ine - t uning _a _large _language _model _involves _adjust ing _the _parameters _of _the _model _to _improve _its _performance _on _a _specific _task _or _down stream _task . _Here _are _some _steps _you _can _follow _to _fine - t une _a _large _language _model : <0x0A> <0x0A> 1 . _Cho ose _the _down stream _task : _Once _you _have _a _large _language _model , ...
(a) Speculative Decoding Example
(b) Tokens Per Autoregressive Iteration
(c) Potential Exploits
Figure 1: (a) In an LLM response with Speculative Decoding (e.g., LADE), tokens are either correctly speculated (blue) and verified in parallel, or mis-speculated (red) and generated via auto-regressive decoding. (b) This pattern can be inferred by measuring the number of tokens generated per iteration, where multiple tokens per iteration indicates correct speculation, and a single token per iteration indicates mis-speculation. (c) Using these patterns, a network-based adversary can fingerprint user queries and learn private user prompts and responses; a malicious user can observe correct predictions and leak out data-stores and hyper-parameters used for predictions.
Consider a response generated by TinyLlama 1.1B Chat V1.0 Zhang et al. (2024a) with the state-of-the-art speculative decoding technique, Lookahead Decoding (LADE) Fu et al. (2024) in Figure 1(a), as an example. The red tokens are mis-speculated, requiring one auto-regressive iteration per token. Whereas, the blue tokens are correctly speculated, with up to $n$ tokens verified in a single iteration, as shown in Figure 1(b). Thus, in LLM applications that stream responses back to the user iteration by iteration, an adversary over the network, similar to recent attacks on LLMs Weiss et al. (2024), can identify network packets corresponding to each iteration and observe the packet size to deduce the number of tokens produced by each iteration and leak the pattern of correct and incorrect speculations dependent on the input.
Exploits. By observing speculation patterns, we demonstrate three possible exploits on speculative-decoding implementations, as shown in Figure 1(c).
First, we demonstrate a query fingerprinting attack by a network-based adversary. By using the speculation-pattern as a fingerprint for an prompt/response pair, the adversary can profile the fingerprints for different prompt/output pairs offline, and then for an unknown user query, measure the speculation-pattern, and compare it with the profiled fingerprints to leak out the user’s private prompt and response. On three different speculative decoding implementations (BiLD, REST, LADE) Fu et al. (2024); He et al. (2024); Kim et al. (2023), we show that one can identify user inputs using fingerprints with accuracy of up to $\sim$100for REST, 92% for LADE, and 95% for BiLD with just 30 training samples. When the prompts to be leaked are just approximately similar to profiled prompts, the accuracy is still within 20% to 40% accuracy (much better than random guessing probability of 2%).
Second, we show that an adversary capable of crafting malicious inputs can leak confidential data or hyper-parameters that may be used for generating predictions in speculative decoding techniques. If predictions are generated via retrieval from a data-store (REST) He et al. (2024), we show that an adversary can craft inputs and monitor which output tokens are correctly predicted, and leak out the data-store at the rate of >25 tokens per second. When speculative decoding generates predictions using self-drafting (LADE) Fu et al. (2024), we show that an adversary can craft malicious inputs that leak prediction hyperparameters like the N-gram size, and the number of prediction candidates generated and verified, which are confidential intellectual property controlling the success rates of a speculative-decoding implementation.
Mitigations. Finally, we discuss potential mitigations against these exploits. We propose transmitting groups of tokens aggregated over several iterations back to the user instead of streaming tokens for each iteration to reduce the observability of fine-grain speculation patterns, and padding packets with a fixed number or random number of tokens to add noise to adversary’s observations to mitigate the fingerprinting attacks. Avoiding usage of private or confidential data for generating predictions also prevents leakage.
In summary, our contributions are as follows:
  1. 1.
    We are the first to observe that variation in the number of tokens generated per autoregressive iteration due to LLM speculative decoding can result in privacy and confidentiality breaches.
  2. 2.
    We demonstrate query fingerprinting attacks that can leak out exact matches of private user queries with >90% accuracy and approximate matches with 20% to 40% accuracy.
  3. 3.
    We also demonstrate attacks that leak confidential IP controlling the performance of the speculative decoding mechanisms, such as data from data-stores used for prediction or hyper-parameters used for predictions.
  4. 4.
    We also propose mitigations including padding packet sizes and aggregating tokens from multiple iterations to limit the potential for these exploits.
2 Background and Motivation
We provide background on speculative decoding in LLMs, and then cover recent side-channel attacks on LLMs.
2.1 Background on Speculative Decoding
Speculative decoding addresses inefficiencies in auto-regressive decoding by allowing multiple token predictions to be verified in parallel, reducing overall latency. It works by generating speculative token guesses, which are then verified by the main model. If correct, multiple tokens are accepted in one step, significantly speeding up the process. If incorrect, the model falls back to standard autoregressive decoding to ensure accuracy.
Speculation Using Smaller Models. Many works propose using smaller, much faster draft models to generate sequences of tokens speculatively Miao et al. (2024); Kim et al. (2023), and then verify a single sequence or a tree of such sequences in parallel using the larger target model. The draft model can be a smaller version of the model family or a pruned larger model Yan et al. (2024). In this paper, we use BiLD Kim et al. (2023), as a representative example of this type. Two key policies—fallback and rollback—ensure the quality of tokens is maintained in BiLD: the fallback policy hands control to the larger model if confidence drops, while rollback corrects tokens exceeding a set distance metric, like cross-entropy loss.
Speculation using Self-Drafting. Another speculative decoding paradigm caches previous output tokens and retrieves them for speculation in future iterations, with the target model verifying these speculations Cai et al. (2024); Fu et al. (2024). Look-Ahead Decoding (LADE) Fu et al. (2024) implements this by maintaining a cache where keys are tokens and values are N-grams previously generated after those tokens. LADE populates its cache using jacobi decoding, selecting random tokens from the input and generating token chains. When a key token reappears in the output stream, LADE uses these N-grams for speculation.
Speculation Using Retrieval. Unlike other paradigms relying on model generations, Retrieval-Based Speculative Decoding (REST) He et al. (2024) speculates by retrieving data from a datastore containing contextually relevant text or code. REST’s datastore contains prefixes and continuations from a corpus, matching input suffixes to generate draft outputs. If there are too many drafts, REST selects high-quality subsets and organizes them in a tree, applying a tree attention mask during decoding with the target model.
2.2 Recent Side-Channel Attacks on LLMs
Recent works have demonstrated side-channel vulnerabilities in LLM inference systems leading to critical privacy risks. A recent attack Weiss et al. (2024) demonstrated a token-length side-channel on streaming LLMs, where network-based attackers can leak a large fraction of private response tokens from an LLM chatbot, by observing the size of encrypted packets in the network traffic. By identifying individual tokens based on packet sizes and leveraging inter-sentence context, this attack reconstructs a significant portion of private responses from platforms like ChatGPT and Microsoft Copilot that stream responses token-by-token. Another attack  Debenedetti et al. (2024) demonstrated a tokenizer side-channel, where a malicious user constructs inputs that translate to different number of tokens based on the tokenizer’s vocabulary, which is used to leak out the vocabulary and any private data in the vocabulary.
2.3 Goal: Speculative Decoding Side-Channel
Speculative execution techniques in processors, deployed for improving performance, have resulted in many potent side-channel attacks capable of leaking information, such as Spectre and Meltdown Kocher et al. (2019); Lipp et al. (2018) and their many variants. Consequently, this work seeks to explore whether the deployment of speculative decoding in LLMs similarly for improving inference performance introduces any new side-channels capable of privacy breaches.
3 Threat Model
Like prior works Weiss et al. (2024), we focus on streaming LLMs, where the user sends requests to an LLM running on the server, and the server sends back the response token-by-token. Here, we study the following attacker models.
  1. 1.
    Network Based Adversary. We assume the attacker is able to observe the network packet size and the time at which they are transmitted, similar to Weiss et al. (2024). The goal of the attacker is to leak private user queries. Although network packets containing LLM response tokens are encrypted, the packet size is still observable. The attacker can also observe the time between each packet to identify which iteration it belongs to. We show this allows an adversary to fingerprint and leak private user queries in Section 4.
  2. 2.
    Malicious User. In this setting, we assume the attacker is a malicious user interacting with the LLM, similar to Debenedetti et al. (2024), with the goal of leaking confidential data used by LLM’s speculative decoding mechanism. Here, the user can craft arbitrary inputs and inspect the output, and also monitor the number of tokens generated per iteration for the output tokens. We show this allows an adversary to leak out any confidential data used for the speculative decoding or its hyperparameters in Section 5.
4 Query Fingerprinting Attack
4.1 Attack Overview
Attack Scenario. We consider a scenario where a user interacts with an AI chatbot using an LLM. The LLM is programmed to answer a predefined set of user queries through the system prompt. This can occur in scenarios such as a patient in a healthcare setting interacting with an LLM to get medical queries answered, or a user in a customer service setting sharing private information with the chatbot before being redirected to a human. The queries are sent from a client device, over a network, to a server that hosts the LLM equipped with speculative decoding. This leads to the threat of a network-based attacker trying to eavesdrop on the queries and responses, as described in Section 3.
Figure 2: Overview of the Query Fingerprinting Attack. In the offline phase, the network-based attacker profiles the trace of number of tokens generated per iteration influenced by speculative decoding, for all possible inputs, and trains a classifier. In the online phase, the attacker extracts the trace of number of tokens generated per iteration for an unknown input, and uses the trained classifier to leak the input.
Goals and Capabilities. Figure 2 shows the overview of our attack. Our attacker over the network seeks to learn the private query asked by the user. While the query and response packets are encrypted, the attacker can intercept the response packets, measure the time intervals between packets, and count the number of tokens per packet or simply use the packet size to approximate the number of tokens.
Mechanism. Since the input set is limited, the attacker fingerprints inputs based on the pattern of output token counts per iteration influenced by speculative decoding. There are two phases to the attack: offline and online. In the offline phase, the attacker profiles the number of output tokens generated per iteration vs iteration-id (we call this a trace) for each input in the input set, and trains a random forest classifier to predict the input based on the trace. In the online phase, the attacker collects the trace for an unknown input and uses the classifier to identify the input. Next, we provide more details about how these fingerprints are obtained, and then describe how the attack works on various speculative decoding techniques (LADE, REST, BiLD).
4.2 Fingerprinting Using Speculation Patterns
Speculative decoding introduces variation in the number of tokens generated per iteration. In iterations with correct speculation, multiple tokens are verified in a single iteration and sent back together in a single packet. However, if the speculation is incorrect, a single token is generated per iteration. The attacker uses this variation to create a fingerprint for a particular output sequence.
As shown in  Figure 2, from each packet containing outputs from a single iteration, the attacker measures the number of tokens generated per iteration. Subsequently, by plotting the token numbers vs iteration ID, it gets the trace for an input/output (prompt-response) tuple.
Figure 3 shows the trace of tokens per iteration for two different prompts with LADE across different runs. From this, we observe two key properties that make these traces suitable for usage to identify and leak the prompt and response:
  1. 1.
    Traces are unique to a prompt and response - different prompts result in different traces of tokens per iteration, as shown in Figure 3(a) and Figure 3(b).
  2. 2.
    Traces are reproducible - repeating the same prompt produces similar patterns in the trace of token counts, as shown in the two iterations in Figure 3(a) or the two iterations in Figure 3(b).
(a) What is a narcissistic personality disorder?
(b) What are the symptoms of cancer?
Figure 3: Trace of tokens generated per iteration vs Iteration-ID, for two runs of different prompts (a) and (b) with LADE. The trace has unique and reproducible patterns, enabling its usage as a fingerprint for leaking prompts.
These observations demonstrate that the number of tokens generated per iteration with speculative decoding is a unique identifier for a prompt/response pair. These patterns can leak the prompt/response via side-channels to an adversary.
However, in existing LLM deployments on servers such as OpenAI’s ChatGPT, the logits get decoded into character strings by the detokenization at the server-side before being streamed to the user in encrypted packets. Because the packets are encrypted, the tokens in each packet cannot be directly inferred. Nevertheless, there is a strong correlation between the packet size and the number of tokens per iteration. Figure 4 shows the trace of token count per iteration for a prompt with LADE, and the corresponding packet size in bytes per iteration match considerably. Thus, for our fingerprinting attack, the adversary can use the trace of packet size per iteration as a proxy for the token count per iteration, given that it is easily observable.
Figure 4: Comparing the trace of token count per iteration and trace of packet size, versus iteration-id for a prompt response with LADE. Strong similarity between the traces allows the attacker to infer speculation patterns, i.e., variation in tokens per iteration, via variation in packet size.
Next, we develop attacks to fingerprint and leak user inputs, based on the traces of packet sizes per iteration.
4.3 Attack Experiment Design
We perform our attacks on an AI Chatbot designed to answer common medical queries from a patient. The user is a patient interacting with the chatbot and the attacker seeks to learn which query the user asks. We conduct the following three fingerprinting attacks:
Experiment 1. The first experiment studies a scenario where the attacker has exact knowledge of the complete set of queries, allowing the attacker to profile the exact queries offline that will be leaked out in the online phase. In this, we use a set of 50 prompts from real world human-LLM interactions acquired from Han et al. (2023). These prompts ask a variety of questions about different common diseases (the list of prompts is provided in the Appendix).
Experiment 2. In this experiment, we choose 50 prompts with commonplace diseases from our dataset that all starts with “What are the symptoms of”. Therefore, all the prompts have a similar structure. We design this experiment to see if our attack can distinguish similar prompts.
The training and testing prompts of both Experiment 1 and 2 are the same. This illustrates the case where the attacker has exact knowledge of the set of predefined prompts.
Experiment 3. The last experiment studies a scenario where the attacker has only approximate knowledge of the set of queries asked by the user. To emulate this, the attacker profiles queries that are semantically similar to the queries asked by the user (e.g., profiled query: “For how long should you take zolpidem?” and user: “How long is zolpidem typically prescribed for?”). We reuse the prompts from Experiment 1 as our training prompts, but for testing, i.e., the user prompts to be leaked, we rephrase the prompts in the training set using ChatGPT-4o by prompting it as "Can you rephrase the following 50 questions while keeping the diseases in each question the same?"
4.4 Methodology
Target Models. We target three recent speculative decoding techniques: LADE Fu et al. (2024) that uses self-drafting for speculation, REST He et al. (2024) that uses retrieval from datastore for speculation, and BiLD Kim et al. (2023) which uses a smaller model for speculation. However, our results are as applicable to other techniques using speculation. We demonstrate results for LADE with TinyLlama 1.1B Chat V1.0 Zhang et al. (2024a), for REST with Vicuna 7B Chiang et al. (2023) as the target model and ShareGPT Aeala (2023) data-store containing 120K conversations used for speculation, and for BiLD with Llama2 7B Touvron et al. (2023) as the target model and TinyLlama 1.1B Zhang et al. (2024a) as the smaller model for speculation. We run LADE and REST on NVIDIA RTX4090 (24GB) and BiLD on NVIDIA A100 (40GB) GPUs with Question-Answer tasks, to simulate a chatbot under attack.
Datasets and Classifier. For each experiment, we use 50 queries of varying complexity, length, and semantics (Appendix lists all the prompts). For our training set, in the offline profiling phase of the attack, we run each query 5 to 30 times and extract the packet size traces (5 to 30 traces per query) and train a Random Forest classifier using them. For the test set in the online phase, we use the corresponding 50 queries and 5 new traces per query, collecting 250 new data points for evaluation, and report accuracy and F1-scores. For our attack, we use scikit-learn’s Random Forest with 150 estimators, max depth of 15, minimum 10 samples per split, minimum 1 sample per leaf, and mean squared error as the loss function.
Accuracy (%) F1 Score
TPQ $\rightarrow$ 5 10 20 30 5 10 20 30
Experiment 1
LADE 53.6% 63.2% 65.5% 68% 0.49 0.59 0.63 0.66
REST 91.2% 97.6% 100% 99.6% 0.91 0.98 1.0 1.0
BiLD 28.4% 41.6% 49.2% 52% 0.27 0.43 0.52 0.55
Experiment 2
LADE 33.2% 53.2% 64.8% 70.8% 0.3 0.51 0.63 0.7
REST 96.4% 97.6% 100% 99.6% 0.96 0.98 1.0 1.0
BiLD 58.8% 68% 75.2% 77.6% 0.54 0.64 0.73 0.76
Experiment 3
LADE 15.2% 16% 13.6% 18.8% 0.13 0.13 0.1 0.16
REST 35.2% 38.4% 40% 39.6% 0.27 0.32 0.34 0.33
BiLD 13.2 % 20% 20.4% 20% 0.11 0.16 0.16 0.16
Table 1: Results of the Query Fingerprinting Attack on three speculative decoding techniques (LADE, REST, BiLD) with temperature of 0.8, with 5 to 30 Traces Per Query (TPQ) used for training the classifier.
4.5 Results
Table 1 shows the results for the Query Fingerprinting Attack on LADE, REST, and BiLD.
Exact Knowledge Scenarios (EK). In both Experiment 1 and 2, the attacker leaks the input prompt with high accuracy: with accuracy of up to 70.8%, 100%, and 77.6% for LADE, REST, and BiLD respectively, compared to random guessing (2% for a set of 50 prompts). In Experiment 1, where we study randomly selected prompts and the set of profiled and leaked prompts are identical, the attack success rates improve as the number of traces per query (TPQ) profiled increases, reaching up to 68% accuracy and an F1 score of 0.66 with 30 TPQ for LADE, reaching 99.6% accuracy with 30 TPQ and an F1 score of 1.0 for REST, and reaching 52% accuracy with 30 TPQ and an F1 score of 0.55 for BiLD. In Experiment 2, where all our prompts have a similar structure, the results are similar, with attack accuracy for LADE, REST, and BiLD reaching 70.8%, 99.6%, and 77.6% respectively. This indicates that our attack success rate is not influenced by the structure of the prompt (input), as long as the response (output) is different.
Approximate Knowledge Scenarios (AK). In Experiment 3, where the leaked prompts are not exactly the same as the profiled prompts, the attack success rates are lower (up to 18.8% for LADE, 20% BiLD, and 40% for REST), but still significantly better than random-guessing (2%). This demonstrates that models using speculative decoding are vulnerable to query fingerprinting attacks, even when the attacker just has approximate knowledge about the queries.
Overall, all of the attacks are more successful on REST than BiLD and LADE. This is because REST demonstrates higher correct speculation rates compared to BiLD and LADE, and therefore has more prompt-dependent variations in the tokens per iteration and the packet sizes, leading to the attack achieving higher accuracy on REST.
4.6 Ablation Studies
Table  2 presents an ablation study studying the attack accuracy as the temperature of our target model varies.
Temperature
0.3 0.6 0.8 1.0
Experiment 1
LADE 91.6% 85.6% 68% 61.2%
REST 100% 100% 99.6% 99.6%
BiLD 74.4% 59.2% 52% 34.8%
Experiment 2
LADE 90.8% 84.4% 70.8% 60.4%
REST 99.6% 100% 99.6% 100%
BiLD 95.2% 88.4% 77.6% 63.6%
Experiment 3
LADE 16.8% 17.6% 18.8% 19.6%
REST 35.2% 42% 39.6% 43.2%
BiLD 17.2% 15.6% 20% 16.4%
Table 2: Accuracy of the Query Fingerprinting Attack on LADE and REST at different temperatures (0.3, 0.6, 0.8, 1.0), with 30 Traces Per Query (TPQ) used to train the classifier.
For LADE and BiLD, as the temperature increases from 0.3 to 1.0, the attack accuracy for both Experiment 1 and Experiment 2 decreases. This is because at lower temperatures, the generations and speculation patterns are more stable and alike. Therefore, the output tokens for the same prompt and the packet size traces are more similar across different runs, making the attack more successful.
For REST, the attack accuracy is relatively stable for both Experiment 1 and 2 despite the increasing temperature. This is likely because the features are so pronounced that it withstands the perturbations caused in the output by the increasing temperature.
The attack accuracy in Experiment 3 has an interesting trend with temperature. Since the attacker only has approximate knowledge of the prompts (train and test set are different), the accuracy of this attack is much lower than Experiment 1 and 2. However, unlike Experiment 1 and 2, for Experiment 3 the attack accuracy has a moderate increase as the temperature of the model increases. This is because, at low temperatures, with limited variation in the training data, the classifier overfits on the training set, and at higher temperatures, this problem does not exist, allowing the attack to reach slightly higher accuracies.
5 Intellectual Property Leakage
Speculative decoding techniques may use carefully selected data or tuned hyper-parameters for achieving the highest correct speculation rates. In this section, we focus on how a malicious user can craft inputs to recover such data or hyper-parameters. We assume the malicious user can examine the output and count the number of tokens streamed back per iteration, as outlined in the threat model in Section 3-2.
5.1 Leaking Datastore Used for Speculation
Techniques like REST He et al. (2024) rely on retrieval from an external datastore to generate speculative tokens. These datastores can be populated with proprietary information to tailor model outputs to specific domains, or data collected from user interactions that is private. Patterns of correct speculations arising from such data can be exploited to leak the sensitive information in such datastores.
5.1.1 Background on REST’s Speculation
REST builds a datastore of phrases in the format of context and continuation pairs, where a preceding context and its following continuation are stored. During generation, REST retrieves the longest context matching with the tokens generated so far, and uses it to construct a speculative draft by selecting top prefixes from the retrieved continuations. This speculative draft is verified using a tree attention mechanism by the target model.
5.1.2 Attack Design
We observe that any tokens that are correctly speculated, i.e., returned as a group of tokens per iteration have to exist in the datastore. We propose an attack where a malicious user crafts inputs and observes which tokens are correctly speculated and all such tokens can thus be leaked from the datastore. With this, the attacker can systematically extract sequences from the datastore.
We study the following attacker strategies for input generation to achieve high token leakage rates:
1. Random Generation: The user prompts the LLM to generate random paragraphs of text.
2. Leveraging Common Words. The LLM is prompted to generate paragraphs containing the most frequently occurring words, taken from a dataset of top 10,000 English words Brants & Franz (2021). This is based on the insight that there is a high chance of phrases in the datastore containing these words.
3. Reusing Leaked Sequences. The LLM is prompted to generate paragraphs that extend phrases leaked out so far, as shown in Figure 5. This approach is based on the insight that the sequence that is already leaked out may be a part of a longer phrase in the datastore. Therefore, with such iterative prompting, additional tokens from the phrase can be leaked out in a focused manner, leading to more unique tokens being leaked with a similar number of queries.
Figure 5: Leakage process where an attacker exploits leaked sequences to extract longer datastore sequences.
Figure 6: Unique sequences leaked from REST datastore as time increases with different attack strategies, averaged over 3 runs per strategy.
5.1.3 Results
We test our attacks on a REST implementation using the ShareGPT datastore of 120K conversations for speculation on NVIDIA RTX 4090. Figure 6 shows the unique sequences leaded by each of the three attack strategies, averaged over three runs per strategy.
Random generation leaks 20,000 unique sequences in 20 minutes, but subsequently the rate slows down leaking only 70,000 sequences in 3 hours. In comparison, prompting with frequently occurring words leaks 31,000 unique sequences in 20 minutes, which increases linearly to 190,000 after 3 hours. Lastly, prompting with leaked sequences results in the most efficient leakage, leaking about 35,000 unique sequences in 20 minutes and about 200,000 after 3 hours, highlighting the amplification potential of feedback. These results show that an attacker can easily leak REST’s datastore when they can observe correct speculation patterns.
5.2 Leakage of Speculation Hyper-parameters
Speculative decoding mechanisms rely on hyper-parameters to control speculation, directly impacting correct speculation rates and performance. By analyzing traces of these patterns, we can reverse-engineer the hyper-parameters used. We demonstrate this attack on LADE Fu et al. (2024).
5.2.1 Background on LADE’s Implementation
LADE generates speculative tokens by caching and reusing prior n-grams generated by the model. LADE has two main parameters: N for n-gram size, and G for guess set size. Its cache is structured as a key-value store, where the key is a token, and the value is a list of G candidate (N-1)-grams following the key token, that are used for speculation. These are replaced in the Least-Recently-Used (LRU) order. During execution, LADE use the last token of the previous iteration or the input to query this cache, and the associated G candidate (N-1)-grams are used as speculation candidates, and the best match with target model generation is accepted.
5.2.2 Leaking Hyper-parameter N in LADE
As each iteration outputs at most N-1 tokens under correct speculation, we expect the maximum number of tokens per iteration to be N-1. Therefore, we prompt LADE with a simple prompt, “Repeat Letter ‘A’ 60 times” that should sustain the maximum correctly speculated tokens per iteration. We observe the maximum number of tokens per iteration with this prompt to deduce N-1, and thus learn N.
Figure 7 shows the number of correct speculated tokens for LADE with three different values of $N$ for this prompt. We see that for each value of N (4, 5, and 6), this prompt has a maximum of 3, 4, and 5 tokens per iteration (N-1), letting us learn the value of N as 4, 5, or 6 in each of these cases. This technique can be extended to learn any value of N used in LADE.
Figure 7: Number of tokens per iteration with LADE for different N, with a prompt that achieves maximum correct speculation (“Repeat Letter ‘A’ 60 times”. The maximum token count per iteration corresponds to value of N-1.)
5.2.3 Leaking Hyper-parameter G in LADE
To leak G, the number of candidate predictions considered for speculation, we force LADE to incur targeted mis-speculations. Here, we prompt the model to repeat a sequence of phrases with the same prefix (“I run”), i.e., “I run to the grocery shop. I run as fast as cars. I run with my best friend. I run from the lecture hall.". If the number of such phrases is less than or equal to G, LADE caches all the N-grams following the key “run”, resulting in it being able to correctly speculate the token that follows “run” in each of these phrases. However, if the number of phrases are greater than G, then with LRU replacement, newer n-grams evict older n-grams for the key “run” from the set of candidates which has a limited capacity of G. This causes the token following “run” to always be mis-speculated.
Figure 8 shows the traces of token generation time for LADE ($G=3$), as the number of phrases ($P$) of the type “I run …” increases from 2 to 4. When $P$ is 2 or 3 ($\leq G$), we notice all the tokens after the initial ones are correctly predicted in the generation (high number of tokens per iteration). But when $P$ is 4 ($>G$), every 7th token, i.e. the token following “run” is mis-speculated (one token per iteration). Thus, we leak the value of G to be 3. This attack can be generalized to any value of parameter G in LADE.
Figure 8: Time Per Token vs Token-ID for LADE with Guess Set (G = 3), using a prompt that asks the LLM to repeat P phrases sharing a common prefix “I run”. Periodic mis-speculations occur for the token after “run”, only when $P>G$, leaking the value of $G$.
6 Mitigation Strategies
In this section, we discuss possible mitigation strategies for our attacks and evaluate their impact on the attack success rate as appropriate.
6.1 Mitigating Query Fingerprinting Attacks
For the query fingerprinting attacks, we propose the following mitigations:
1. Padding Token Packets: Padding the network packet size with additional bytes can mask the actual number of tokens per iteration and conceal the differences between correct and mis-speculations to prevent the attack.
The padding can be added in two ways:
  1. 1.
    Constant Length Padding. Every packet can be padded to a constant size. This size can be selected based on the maximum number of tokens per iteration, multiplied by the maximum bytes per token. We pick this size as 1024 bytes, without loss of generality.
  2. 2.
    Variable Length Padding. Every packet can be padded with a random number of bytes selected from a uniform distribution of $\epsilon\sim Unif(0,D)$ where D is a pre-defined upper bound of padding and $\epsilon$ is an integer.
Table 3: Accuracy after padding each packet to 1024 bytes (attack: 20 traces per prompt; model temperature: 0.8)
No Padding Padding to 1024 Bytes
Experiment 1
LADE 65.6% 3.2%
REST 100% 2%
BiLD 49.2% 2%
Experiment 2
LADE 64.8% 2.4%
REST 100% 2%
BiLD 75.2% 2%
Experiment 3
LADE 13.6% 4%
REST 40% 2%
BiLD 20.4% 2%
Constant Length Padding. Table 3 shows the accuracy of the attack after padding every response packet to a constant size of 1024 Bytes. This completely mitigates the side channel, reducing the attack accuracy to close to random guessing (2%). However, this comes at the cost of considerable communication overhead and inefficiency, due to an increase in packet size by almost 230$\times$, as majority of the packet consists of random padding bytes.
Table 4: Attack accuracy after padding each packets with random bytes from uniform distribution, $\epsilon\sim Unif(0,D)$, where D varies from 6 to 48. (attack: 20 traces per prompt; model temperature: 0.8)
Accuracy vs Max Padding Size (D) Overhead vs D
D $\rightarrow$ 0 6 12 24 48 6 12 24 48
Experiment 1
LADE 65.6% 25.6% 13.6% 8.4% 3.6% 1.5$\times$ 2$\times$ 3.1$\times$ 5.1$\times$
REST 100% 90.4% 75.6% 48% 27.2% 1.9$\times$ 2.8$\times$ 4.6$\times$ 8.2$\times$
BiLD 49.2% 26.8% 11.6% 5.2% 4.8% 1.7$\times$ 2.3$\times$ 3.6$\times$ 6.3$\times$
Experiment 2
LADE 64.8% 19.2% 10.8% 5.6% 4.4% 1.6$\times$ 2.1$\times$ 3.2$\times$ 5.4$\times$
REST 100% 95.2% 76.8% 52% 34.4% 2$\times$ 3$\times$ 4.8$\times$ 8.7$\times$
BiLD 75.2% 37.2% 21.6% 8.4% 5.6% 1.7$\times$ 2.4$\times$ 3.7$\times$ 6.5$\times$
Experiment 3
LADE 13.6% 6.8% 6% 3.6% 3.2% 1.5$\times$ 2$\times$ 3$\times$ 5$\times$
REST 40% 29.6% 20.8% 14.8% 8.4% 1.9$\times$ 2.8$\times$ 4.6$\times$ 8.1$\times$
BiLD 20.4% 7.6% 6.8% 6% 3.6% 1.7$\times$ 2.3$\times$ 3.6$\times$ 6.2$\times$
Variable Length Padding. Table 4 shows the attack accuracy after padding responses with a variable number of bytes, $\epsilon\sim Unif(0,D)$, where D varies from 6 to 48. As D increases to 48, the attack accuracy for LADE and BiLD drops close to random guess (2%) for all experiments. The attack accuracy for REST also shows a considerable drop from 100% without padding to 27% and 34% with padding of up to 48 bytes for Experiments 1 and 2 respectively. Similarly, for Experiment 3, attack accuracy drops from 40% to 8% with D=48. These reductions in accuracy come with a 5$\times$ to 8$\times$ increase in the packet size for D=48. Thus, variable length padding provides a better trade-off of reduction in attack accuracy with limited increase in packet size compared to constant length padding.
2. Reduce Granularity of Attacker Observation: An alternative approach is to aggregate output tokens over multiple iterations at the server and return them to the client all together. This can limit observability of fine-grain speculation patterns from the packet sizes for the attacker, and reduce the effectiveness of fingerprinting attacks.
Table 5 shows the results for the server aggregating output tokens over 3 to 20 iterations. The accuracy of the exact knowledge attacks (Experiment 1 and 2) for LADE, REST, and BiLD drops to a minimum of 9.2%, 52%, and 11.2% respectively after aggregating the tokens over up to 20 iterations down from 65%, 100%, and 49.2% respectively without aggregation; accuracy of the approximate knowledge attacks (Experiment 3) decreases to a minimum of 5.2%, 17.6%, and 4% respectively, down from 14%, 40%, and 20% respectively.
Thus, reducing granularity is an effective mitigation against query fingerprinting attacks. One downside of this approach is that the user waits longer between updates to the output as the number of iterations the tokens are aggregated over increases. This limits the number of iterations tokens can be aggregated over without hindering system responsiveness.
Table 5: Accuracy of fingerprinting attacks (20 traces per prompt and temperature of 0.8) with aggregation of tokens across 3 to 20 iterations.
Aggregation Granularity (Number of Iterations)
1 3 5 10 20
Experiment 1
LADE 65.6% 53.2% 40% 24.4% 10.8%
REST 100% 96.8% 89.2% 80.4% 52%
BiLD 49.2% 34% 30% 19.2% 11.2%
Experiment 2
LADE 64.8% 45.2% 28.8% 16.4% 9.2%
REST 100% 97.6% 92.8% 75.6% 52%
BiLD 75.2% 57.6% 41.6% 30% 16.8%
Experiment 3
LADE 13.6% 13.2% 14.4% 9.2% 5.2%
REST 40% 33.2% 28.8% 20.4% 17.6%
BiLD 20.4% 15.6% 10% 10% 4%
6.2 Mitigating Datastore or Hyper-parameter Leaks
For preventing confidential data leakage from datastore or parameter leakage attacks, in addition to the two mitigations discussed previously, we can consider one other mitigation:
3. Using Only Public Data for Speculative Decoding: Any data-store used for speculation (e.g., in REST) should be constructed only with public data, and any private data or personal identifiable information (e.g., usernames, address, etc.) should be scrubbed from it before it is used. This minimizes the risk of sensitive data leakage through any speculation patterns that may be observable despite padding or token aggregation. Similarly, parameters used for speculation (such as N and G in LADE) should be assumed to be public as they can be easily leaked out.
7 Related Work
In this section, we describe other speculative decoding techniques that may be exploitable and related LLM performance optimizations that may lead to side-channel leakage.
7.1 LLM Speculative Decoding Techniques
Speculative Decoding with Tree Attention. SpecInfer Miao et al. (2024) and SpecTr Sun et al. (2023) observe that when a token early in the speculation draft is rejected, the entire draft gets discarded, resulting in wasted computation. Instead, these works propose multiple drafts per iteration organized in a tree structure to provide alternatives for speculative tokens at each position, increasing the likelihood of tokens advancing to the next step and yielding greater speedup by increasing correct speculation rates. These drafts are organized in a tree structure with a custom tree attention mask to prevent inter-tree interactions and verified in parallel. Such works (similar to REST He et al. (2024) which also uses tree attention) increase the correct speculation rate of tokens and introduce more input-dependent speculation patterns in the output. Hence these are likely to be more vulnerable to our query fingerprinting attacks, similar to REST, which is highly susceptible to our attack, as we show in this paper.
Speculative Decoding using Target Models. Other works reuse the target model itself for speculative decoding by utilizing the feature layer (second-to-top layer). Medusa Cai et al. (2024) trains multiple extra heads, each responsible for a token position in the draft sequence, to generate speculative tokens. EAGLE and EAGLE-2 Li et al. (2024a; b) propose a method to perform auto-regressive decoding on the feature layers itself to generate speculative tokens. These techniques are conceptually no different from using an alternative smaller model for speculation like BiLD. Hence, they are equally susceptible to query fingerprinting attacks.
7.2 Performance Optimizations in LLMs
Other performance optimization techniques in LLMs that can introduce timing variation, may also be susceptible to leakage via timing side-channels.
Adaptive KV-Cache Management. The KV-cache is the bottleneck for token generation latency for the memory-bound LLM inference Pope et al. (2023). Recent works, like H2O Zhang et al. (2024b) that performs input-dependent cache eviction or TriForceSun et al. (2024) that performs retrieval-based KV-cache selection for drafting may have input-dependent timing variations. These can introduce timing side-channels allowing query fingerprinting, similar to our attacks. However, as such attacks rely on timing, they are likely to be more susceptible to noise.
Prompt Caching A significant portion of prompts sent to LLMs are duplicates, including system prompts, repeated inputs, and others. Prompt caching, adopted by major LLM providers such as Anthropic and OpenAI Anthropic (2024); OpenAI (2024b); Gim et al. (2024), precomputes attention states for these repetitive prompt components to enable reuse, and substantially reducing latency for the first token when there is reuse. Such techniques can leak the nature of the query based on the timing. Other works cache the responses for prompts with similar semantic meanings Bang (2023). Recent attacks Song et al. (2024) demonstrate query fingerprinting attacks similar to ours, when such prompt caches are shared between users. In contrast, our attacks are effective even without any explicit sharing of caches.
8 Conclusion
This paper reveals significant privacy risks associated with speculative decoding in Large Language Models. Across multiple speculative decoding techniques, we demonstrate that attacks can exploit speculation patterns to infer user inputs with $>90$% accuracy via fingerprinting attacks and also leak data from datastores used for predictions. These leakages highlight the careful consideration needed when deploying speculative decoding techniques to ensure that performance does not come at the cost of reduced privacy.
Appendix A Prompts for Fingerprinting Attacks
A.1 Training and testing prompts for Experiment 1
What to expect if I have Allergy (Outlook/Prognosis)?
What causes Cardiogenic shock?
What to expect if I have Congenital heart block (Outlook/Prognosis)?
What are the symptoms of Congestive heart failure?
What are the symptoms of Coronary heart disease?
What to expect if I have Dengue fever (Outlook/Prognosis)?
What causes Stroke?
Who is at highest risk for Heart attack?
What to expect if I have Heart murmur (Outlook/Prognosis)?
When to seek urgent medical care when I have Metabolic syndrome?
What are the symptoms of Palpitation?
What causes Radiation injury?
What are the symptoms of shock?
Who is at highest risk for Back pain?
What are the symptoms of Athlete’s foot?
What causes Boil?
When to seek urgent medical care when I have Burns?
What causes Corns & calluses?
Who is at highest risk for Fleas?
When to seek urgent medical care when I have Hand-foot-and-mouth disease?
When to seek urgent medical care when I have Hives?
When to seek urgent medical care when I have Liver spots?
What to expect if I have Measles (Outlook/Prognosis)?
When to seek urgent medical care when I have Mumps?
What are the symptoms of a Rash?
When to seek urgent medical care when I have Ringworm?
What are the symptoms of Scarlet fever?
What causes Shingles?
What are the symptoms of Sunburn?
When to seek urgent medical care when I have Wart?
What are the symptoms of Common cold?
Who is at highest risk for AIDS ?
What are the causes of Alcoholic liver disease?
What to expect if I have Anal cancer (Outlook/Prognosis)?
When to seek urgent medical care when I have Anal fissure?
When to seek urgent medical care when I have Anemia?
When to seek urgent medical care when I have Antisocial personality disorder?
How many children have Autism?
What are the symptoms of Avian influenza?
When to seek urgent medical care when I have Avoidant personality disorder?
What causes cataract?
Who is at highest risk for Chickenpox?
What causes Chronic fatigue syndrome?
When to seek urgent medical care when I have Depression?
What to expect if I have Color blindness?
How many children have Dependent personality disorder?
When to seek urgent medical care when I have Adolescent depression?
Who is at highest risk for Diabetes?
What to expect if I have Ebola?
What causes fatty liver?
A.2 Training and testing prompts for Experiment 2
What are the symptoms of Allergy?
What are the symptoms of Cardiogenic shock?
What are the symptoms of heart block?
What are the symptoms of Congestive heart failure?
What are the symptoms of Coronary heart disease?
What are the symptoms of Dengue fever?
What are the symptoms of stroke?
What are the symptoms of a Heart attack?
What are the symptoms of Heart murmur?
What are the symptoms of Metabolic syndrome?
What are the symptoms of Palpitation?
What are the symptoms of Radiation injury?
What are the symptoms of shock?
What are the symptoms of Back pain?
What are the symptoms of Athlete’s foot?
What are the symptoms of Boil?
What are the symptoms of Burns?
What are the symptoms of Corns & calluses?
What are the symptoms of Fleas?
What are the symptoms of Hand-foot-and-mouth disease?
What are the symptoms of Hives?
What are the symptoms of Liver spots?
What are the symptoms of Measles?
What are the symptoms of Mumps?
What are the symptoms of a Rash?
What are the symptoms of Ringworm?
What are the symptoms of Scarlet fever?
What are the symptoms of Shingles?
What are the symptoms of Sunburn?
What are the symptoms of Warts?
What are the symptoms of Common cold?
What are the symptoms of AIDS?
What are the symptoms of Alcoholic liver disease?
What are the symptoms of Anal cancer?
What are the symptoms of an Anal fissure?
What are the symptoms of Anemia?
What are the symptoms of Antisocial personality disorder?
What are the symptoms of Autism?
What are the symptoms of Avian influenza?
What are the symptoms of Avoidant personality disorder?
What are the symptoms of a cataract?
What are the symptoms of Chickenpox?
What are the symptoms of Chronic fatigue syndrome?
What are the symptoms of Depression?
What are the symptoms of Color blindness?
What are the symptoms of Dependent personality disorder?
What are the symptoms of Adolescent depression?
What are the symptoms of Diabetes?
What are the symptoms of Ebola?
What are the symptoms of fatty liver?
A.3 Training prompts for Experiment 3
Same as Section A.1
A.4 Testing prompts for Experiment 3
What is the outlook for someone with an allergy?
What leads to cardiogenic shock?
How might congenital heart block affect me in the long term?
How does congestive heart failure present itself?
How can coronary heart disease symptoms be identified?
What is the prognosis for someone with dengue fever?
What are the main causes of a stroke?
Who faces the greatest risk for a heart attack?
What is the long-term prognosis for someone with a heart murmur?
When is it critical to get help for metabolic syndrome?
How can I tell if I’m experiencing palpitations?
What factors contribute to radiation injury?
What are the warning signs of shock?
Who is most susceptible to back pain?
What should I look out for if I suspect athlete’s foot?
What are the main causes of a boil?
When are burns severe enough to seek emergency medical care?
Why do corns and calluses form?
Who is most at risk of getting fleas?
When should I go to the doctor for hand-foot-and-mouth disease?
When is hives a sign to seek urgent medical attention?
When should liver spots be evaluated by a doctor?
What can I expect if I have measles?
When should I seek medical help for mumps?
How do I know if I have a rash?
When should ringworm be treated urgently?
What are the common symptoms of scarlet fever?
What is the underlying cause of shingles?
How can I recognize sunburn symptoms?
When should a wart be looked at by a doctor?
What are the signs of the common cold?
Who is most at risk of contracting HIV/AIDS?
What factors contribute to alcoholic liver disease?
What is the outlook for someone diagnosed with anal cancer?
When is an anal fissure an emergency?
When should anemia symptoms prompt immediate medical care?
When is antisocial personality disorder considered an urgent issue?
How common is autism in children?
What symptoms are typical of avian influenza?
When should avoidant personality disorder be addressed by a professional?
What are the main causes of cataracts?
Who is most vulnerable to contracting chickenpox?
What are the underlying causes of chronic fatigue syndrome?
When should I seek emergency medical help if I have depression?
What are the typical symptoms or outcomes if I have color blindness?
How common is dependent personality disorder in children?
When is urgent medical attention needed for adolescent depression?
Who faces the highest risk of developing diabetes?
What should I expect if I have Ebola?
What leads to the development of fatty liver?

### v3

@@
-
+When Speculation Spills Secrets: Side Channels via Speculative Decoding in LLMs
 Abstract
-Speculative decoding in large language models (LLMs) accelerates token generation by speculatively predicting multiple tokens cheaply and verifying them in parallel, and has been widely deployed.
-In this paper, we provide the first study demonstrating the privacy risks of speculative decoding.
-We observe that input-dependent patterns of correct and incorrect predictions can be leaked out to an adversary monitoring token generation times and packet sizes, leading to privacy breaches.
-By observing the pattern of correctly and incorrectly speculated tokens, we show that a malicious adversary can fingerprint queries and learn private user inputs with more than 90% accuracy across three different speculative decoding techniques – REST (almost 100% accuracy), LADE (up to 92% accuracy), and BiLD (up to 95% accuracy).
-We show that an adversary can also leak out confidential intellectual property used to design these techniques, such as data from data-stores used for prediction (in REST) at a rate of more than 25 tokens per second, or even hyper-parameters used for prediction (in LADE).
-We also discuss mitigation strategies, such as aggregating tokens across multiple iterations and padding packets with additional bytes, to avoid such privacy or confidentiality breaches.
+Deployed large language models (LLMs) often rely on speculative decoding, a technique that generates and verifies multiple candidate tokens in parallel, to improve throughput and latency.
+In this work, we reveal a new side-channel whereby input-dependent patterns of correct and incorrect speculations can be inferred by monitoring per-iteration token counts or packet sizes.
+We demonstrate that an adversary observing these patterns can fingerprint user queries with ¿90% accuracy across four speculative-decoding schemes, REST ($\sim$100%), LADE (up to 92%), BiLD (up to 95%), and EAGLE (up to 77.6%) and leak confidential datastore contents used for prediction at rates exceeding 25 tokens/sec.
+We evaluate the side-channel attacks in both research prototypes as well as the production-grade vLLM serving framework.
+To defend against these, we propose and evaluate a suite of mitigations, including packet padding and iteration-wise token aggregation.
 1 Introduction
-Large Language Models (LLMs) have transformed natural language processing (NLP), enabling machines to generate and understand human language at an unprecedented scale Vaswani et al.
+Large Language Models (LLMs) have transformed natural language processing (NLP), allowing machines to generate and understand human language at an unprecedented scale Vaswani et al.
 (2017); Devlin et al.
@@
 LLMs typically generate text using auto-regressive decoding Touvron et al.
-(2023); OpenAI (2024a), where the generation happens serially and each token depends on all the previous ones.
+(2023); OpenAI (2024), where the generation happens serially and each token depends on all the previous ones.
 Unfortunately, this serial process causes a significant bottleneck in the LLM response latency Miao et al.
@@
 (2024) and under-utilizes the available hardware-level parallelism, limiting token generation throughput and latency.
-Speculative decoding Leviathan et al.
-(2023); Chen et al.
-(2023); Miao et al.
-(2024); Spector & Re (2023) addresses this problem without impacting model accuracy.
-It uses smaller models or heuristics such as retrieval or self-drafting to inexpensively generate tokens speculatively, which the larger target model verifies in parallel in a single iteration.
-By tuning the heuristics to maintain a high rate of correct speculations, such techniques provide 2x to 5x speedups in inference latency and throughput Xia et al.
-(2024).
-However, the adoption of speculative techniques is not without risks.
-Speculative execution in CPUs  Burton (1985); Hennessy & Patterson (2012), which inspired speculative decoding, has led to considerable security vulnerabilities in processors, such as Spectre Kocher et al.
-(2019), Meltdown Lipp et al.
-(2018), and their numerous variants.
-These side-channel vulnerabilities exploit the side-effects of processor mis-speculations, such as timing variations introduced by mis-speculations, to leak private and confidential data that become accessible to a processor during mis-speculations.
+[…]
+However, the adoption of speculation techniques is not without risks.
+Speculative execution in CPUs  Burton (1985); Hennessy & Patterson (2012), which inspired speculative decoding, has led to security vulnerabilities in processors, such as Spectre Kocher et al.
+(2019) and their variants, which exploit side-channels, i.e.
+timing variations due to mis-speculations that leak secret data accessed during mis-speculations.
 This raises the question, does speculative decoding in LLMs also introduce new risks to privacy?
-In this paper, we provide the first study of the privacy and confidentiality risks of speculative decoding in LLMs, including leakage of private user inputs or confidential data.
+In this paper, we provide a study of the privacy risks of speculative decoding in LLMs, including leakage of private user inputs.
 Problem.
@@
 Q.
-How to fine-tune a large language model?
+How to fine-tune a LLM?
 A.
 F ine - t uning _a _large _language _model _involves _adjust ing _the _parameters _of _the _model _to _improve _its _performance _on _a _specific _task _or _down stream _task .
-_Here _are _some _steps _you _can _follow _to _fine - t une _a _large _language _model : <0x0A> <0x0A> 1 .
-_Cho ose _the _down stream _task : _Once _you _have _a _large _language _model , ...
-(a) Speculative Decoding Example
-(b) Tokens Per Autoregressive Iteration
+_Here _are _some _steps _you _can _follow _to _fine - t une _a _large _language
+(a) Spec Decoding Example
+(b) Tokens Per Iteration
 (c) Potential Exploits
@@
 (c) Using these patterns, a network-based adversary can fingerprint user queries and learn private user prompts and responses; a malicious user can observe correct predictions and leak out data-stores and hyper-parameters used for predictions.
-Consider a response generated by TinyLlama 1.1B Chat V1.0 Zhang et al.
-(2024a) with the state-of-the-art speculative decoding technique, Lookahead Decoding (LADE) Fu et al.
-(2024) in Figure 1(a), as an example.
-The red tokens are mis-speculated, requiring one auto-regressive iteration per token.
-Whereas, the blue tokens are correctly speculated, with up to $n$ tokens verified in a single iteration, as shown in Figure 1(b).
-Thus, in LLM applications that stream responses back to the user iteration by iteration, an adversary over the network, similar to recent attacks on LLMs Weiss et al.
-(2024), can identify network packets corresponding to each iteration and observe the packet size to deduce the number of tokens produced by each iteration and leak the pattern of correct and incorrect speculations dependent on the input.
-Exploits. By observing speculation patterns, we demonstrate three possible exploits on speculative-decoding implementations, as shown in Figure 1(c).
-First, we demonstrate a query fingerprinting attack by a network-based adversary.
-By using the speculation-pattern as a fingerprint for an prompt/response pair, the adversary can profile the fingerprints for different prompt/output pairs offline, and then for an unknown user query, measure the speculation-pattern, and compare it with the profiled fingerprints to leak out the user’s private prompt and response.
-On three different speculative decoding implementations (BiLD, REST, LADE) Fu et al.
+Figure 1(a) shows a response by TinyLlama 1.1B Chat  Zhang et al.
+(2024a) with a speculative decoding technique, Lookahead Decoding (LADE) Fu et al.
+(2024), with correct (blue) and mis-speculated (red) tokens illustrated.
+On correct speculation, up to $n$ tokens are verified per iteration, as shown in Figure 1(b), while on mis-speculation, only one token is generated per iteration.
+In LLM applications streaming responses to the user, token by token, a network-based adversary, such as a malicious ISP or a compromised router Weiss et al.
+(2024), can observe network packet sizes corresponding to each iteration to deduce the number of tokens per iteration and learn the input-dependent pattern of correct and incorrect speculations.
+We target medical chatbots (e.g., Hippocratic AI Shah (2025)), where users interact by asking questions about symptoms or diseases, as this is a privacy-sensitive application where any information leakage (e.g., disease or symptoms of users) breaks privacy laws.
+Exploits.
+We demonstrate privacy breaches based on speculation patterns, as shown in Figure 1(c).
+First, we show a query fingerprinting attack (e.g., leaking diseases or symptoms in medical chatbots) by a network adversary.
+By profiling speculation pattern fingerprints for prompt-response pairs offline, an adversary can then use compare the speculation pattern for an unknown query with the fingerprints to recover private user prompts.
+Across speculative decoding schemes (BiLD, REST, LADE) Fu et al.
 (2024); He et al.
 (2024); Kim et al.
-(2023), we show that one can identify user inputs using fingerprints with accuracy of up to $\sim$100for REST, 92% for LADE, and 95% for BiLD with just 30 training samples.
-When the prompts to be leaked are just approximately similar to profiled prompts, the accuracy is still within 20% to 40% accuracy (much better than random guessing probability of 2%).
-Second, we show that an adversary capable of crafting malicious inputs can leak confidential data or hyper-parameters that may be used for generating predictions in speculative decoding techniques.
-If predictions are generated via retrieval from a data-store (REST) He et al.
-(2024), we show that an adversary can craft inputs and monitor which output tokens are correctly predicted, and leak out the data-store at the rate of >25 tokens per second.
-When speculative decoding generates predictions using self-drafting (LADE) Fu et al.
-(2024), we show that an adversary can craft malicious inputs that leak prediction hyperparameters like the N-gram size, and the number of prediction candidates generated and verified, which are confidential intellectual property controlling the success rates of a speculative-decoding implementation.
-Mitigations.
-Finally, we discuss potential mitigations against these exploits.
-We propose transmitting groups of tokens aggregated over several iterations back to the user instead of streaming tokens for each iteration to reduce the observability of fine-grain speculation patterns, and padding packets with a fixed number or random number of tokens to add noise to adversary’s observations to mitigate the fingerprinting attacks.
-Avoiding usage of private or confidential data for generating predictions also prevents leakage.
+(2023), our prompt identification attacks reaches accuracies of $\sim$100% for REST, 92% for LADE, and 95% for BiLD with only 30 training samples.
+On a remote vLLM inference server Kwon et al.
+(2023), which supports speculation with EAGLE Li et al.
+(2024a), we similarly observe a high accuracy of 77.6%.
+We further show that when the actual prompts are unavailable, the adversary can train effective fingerprints using publicly available proxy datasets; e.g., by querying a medical chatbot with the top 50 most common diseases as stand-ins for real prompts.
+In this case, the attack still achieves 20–40% accuracy in recovering the disease or symptoms from user prompts, far exceeding random guessing (2%).
+Lastly, we show how adversaries crafting malicious inputs can also leak confidential data such as data-store contents used for predictions with REST He et al.
+(2024) at a rate of more than 25 tokens/second by observing correctly predicted tokens.
+[…]
 In summary, our contributions are as follows:
   1. 1.
-    We are the first to observe that variation in the number of tokens generated per autoregressive iteration due to LLM speculative decoding can result in privacy and confidentiality breaches.
+    We observe that variation in the number of tokens generated per autoregressive iteration due to LLM speculative decoding can result in privacy breaches.
   2. 2.
-    We demonstrate query fingerprinting attacks that can leak out exact matches of private user queries with >90% accuracy and approximate matches with 20% to 40% accuracy.
+    We demonstrate query fingerprinting attacks that can leak out exact matches of private user queries with >>90% accuracy and approximate matches with 20% to 40% accuracy.
   3. 3.
-    We also demonstrate attacks that leak confidential IP controlling the performance of the speculative decoding mechanisms, such as data from data-stores used for prediction or hyper-parameters used for predictions.
+    We also demonstrate attacks that leak confidential IP controlling the performance of the speculative decoding mechanisms, such as data from data-stores used for prediction.
   4. 4.
-    We also propose mitigations including padding packet sizes and aggregating tokens from multiple iterations to limit the potential for these exploits.
-2 Background and Motivation
-We provide background on speculative decoding in LLMs, and then cover recent side-channel attacks on LLMs.
-2.1 Background on Speculative Decoding
-Speculative decoding addresses inefficiencies in auto-regressive decoding by allowing multiple token predictions to be verified in parallel, reducing overall latency.
-It works by generating speculative token guesses, which are then verified by the main model.
-If correct, multiple tokens are accepted in one step, significantly speeding up the process.
-If incorrect, the model falls back to standard autoregressive decoding to ensure accuracy.
-Speculation Using Smaller Models.
-Many works propose using smaller, much faster draft models to generate sequences of tokens speculatively Miao et al.
+    We propose mitigations including padding packet sizes and aggregating tokens from multiple iterations to limit the potential for these exploits.
+2 Related Work
+2.1 Prior Speculative Decoding Techniques
+Speculative decoding improves the efficiency of auto-regressive decoding in LLMs by verifying multiple token predictions in parallel, reducing latency.
+These works generate tokens using a smaller model or other heuristics as predictions, which are verified by the target model.
+Speculation with Smaller Draft Models.
+Prior works propose using smaller, faster draft models to generate sequences of tokens speculatively Miao et al.
 (2024); Kim et al.
@@
 (2025).
-In this paper, we use BiLD Kim et al.
-(2023), as a representative example of this type.
-Two key policies—fallback and rollback—ensure the quality of tokens is maintained in BiLD: the fallback policy hands control to the larger model if confidence drops, while rollback corrects tokens exceeding a set distance metric, like cross-entropy loss.
-Speculation using Self-Drafting.
-Another speculative decoding paradigm caches previous output tokens and retrieves them for speculation in future iterations, with the target model verifying these speculations Cai et al.
-(2024); Fu et al.
+In this paper, we demonstrate our attacks on BiLD Kim et al.
+(2023), as a example of this type of speculation.
+Speculation via Self-Drafting.
+Recent works reuse the target model for producing predictions, eliminating the need for a separate draft model.
+Look-Ahead Decoding (LADE) Fu et al.
+(2024) caches previous output tokens and uses them to speculate future tokens via a key-value store populated with token sequence N-grams.
+These are populated using Jacobi decoding, by selecting random tokens from the input and generating successive chain of tokens.
+When this key token reappears in the output stream, LADE uses the stored N-grams for speculation.
+Medusa Cai et al.
+(2024) trains extra heads in the target model, each responsible for a token position in the draft sequence, for speculative tokens.
+EAGLE and EAGLE-2 Li et al.
+(2024a; b) perform auto-regressive decoding directly on the feature layers itself to generate speculative tokens, and are integrated in the vLLM inference serving platform.
+Our attacks target LADE (locally) and EAGLE (on a remote vLLM server) as representative examples, but our techniques are broadly applicable to other techniques as well.
+Speculation Using Retrieval.
+REST He et al.
+(2024) generates predictions by retrieving data from a datastore containing relevant text or code, organized as a set of prefixes and continuations that can be potentially used as speculative tokens.
+If there are multiple candidates, REST organizes them in a tree, applying a tree attention mask during decoding with the target model.
+Tree-Based Verification.
+SpecInfer Miao et al.
+(2024), SpecTr Sun et al.
+(2023) and REST He et al.
+(2024) generate multiple drafts of speculation per iteration, organized in a tree structure, to increase the length of successful verification for higher speedup.
+These drafts are organized with a custom tree attention mask to prevent inter-tree interactions and verified in parallel.
+In this work, we demonstrate our attacks on REST He et al.
+(2024), as a representative example of such techniques.
+2.2 Side-Channel Attacks on LLMs
+Debenedetti et al.
+(2024) pioneered side-channel attacks on LLMs.
+Their attack exploits variation in output token counts due to the tokenizer, to leak uncommon strings in the tokenizer vocabulary.
+Our work instead exploits variation in token count per iteration due to speculative decoding.
+Weiss et al.
+(2024) identified a token-length side channel in streaming LLMs, where network attackers can infer token values by analyzing encrypted packet sizes.
+This attack reconstructs around 27% of the LLM outputs using the character length of each token.
+In our work, we similarly focus on a network-attacker observing the encrypted packet sizes, and leverage the variation in packet sizes based on speculative-decoding to identify user inputs.
+Moreover, we show mitigating the token-length side channel by padding tokens to the maximum token-size still results in packet size variations due to speculative decoding varying the output token counts per packet.
+We show that this can be used to fingerprint prompts and responses in our attacks with greater than 50% accuracy, even after the token-length side-channel is mitigated, in Appendix D.
+Carlini & Nasr (2024) demonstrated timing attacks on LLM inference optimizations in commercial models like GPT-4 and Claude.
+Their attacks rely on timing variations due to these optimizations, i.e.
+variations in packet inter-arrival times, which is considerably susceptible to noise due to network congestion wand server loads.
+In comparison, our attacks exploit packet size variations due to inference optimizations, which are inherently more robust and independent of server or network load variations.
+We show the resilience of our attacks compared to this prior work in Section 4.7.
+Song et al.
+(2024) and  Zheng et al.
+(2024) demonstrated timing side-channel attacks exploiting shared key-value (KV) and semantic caches to leak sensitive LLM inputs.
+Wiretapping LLMs Soleimani et al.
+(2025) also reveals private attributes in LLM-based medical and financial applications.
+3 Threat Model
+Like prior works Weiss et al. (2024), we focus on streaming LLMs, where the LLM server sends back responses to the user token-by-token. Here, we study the following attacker models.
+Network Attacker.
+We assume a malicious ISP or a compromised router that can observe network packet sizes, similar to Weiss et al.
 (2024).
-Look-Ahead Decoding (LADE) Fu et al.
-(2024) implements this by maintaining a cache where keys are tokens and values are N-grams previously generated after those tokens.
-LADE populates its cache using jacobi decoding, selecting random tokens from the input and generating token chains.
-When a key token reappears in the output stream, LADE uses these N-grams for speculation.
-Speculation Using Retrieval.
-Unlike other paradigms relying on model generations, Retrieval-Based Speculative Decoding (REST) He et al.
-(2024) speculates by retrieving data from a datastore containing contextually relevant text or code.
-REST’s datastore contains prefixes and continuations from a corpus, matching input suffixes to generate draft outputs.
-If there are too many drafts, REST selects high-quality subsets and organizes them in a tree, applying a tree attention mask during decoding with the target model.
-2.2 Recent Side-Channel Attacks on LLMs
-Recent works have demonstrated side-channel vulnerabilities in LLM inference systems leading to critical privacy risks.
-A recent attack Weiss et al.
-(2024) demonstrated a token-length side-channel on streaming LLMs, where network-based attackers can leak a large fraction of private response tokens from an LLM chatbot, by observing the size of encrypted packets in the network traffic.
-By identifying individual tokens based on packet sizes and leveraging inter-sentence context, this attack reconstructs a significant portion of private responses from platforms like ChatGPT and Microsoft Copilot that stream responses token-by-token.
-Another attack  Debenedetti et al.
-(2024) demonstrated a tokenizer side-channel, where a malicious user constructs inputs that translate to different number of tokens based on the tokenizer’s vocabulary, which is used to leak out the vocabulary and any private data in the vocabulary.
-2.3 Goal: Speculative Decoding Side-Channel
-Speculative execution techniques in processors, deployed for improving performance, have resulted in many potent side-channel attacks capable of leaking information, such as Spectre and Meltdown Kocher et al.
-(2019); Lipp et al.
-(2018) and their many variants.
-Consequently, this work seeks to explore whether the deployment of speculative decoding in LLMs similarly for improving inference performance introduces any new side-channels capable of privacy breaches.
-3 Threat Model
-Like prior works Weiss et al.
-(2024), we focus on streaming LLMs, where the user sends requests to an LLM running on the server, and the server sends back the response token-by-token.
-Here, we study the following attacker models.
-  1. 1.
-    Network Based Adversary.
-We assume the attacker is able to observe the network packet size and the time at which they are transmitted, similar to Weiss et al.
+The attacker’s goal is to leak private user queries.
+Although LLM response packets are encrypted, the packet size is still observable.
+The attacker can also observe the time between each packet to identify the iteration ID.
+We show how an adversary can fingerprint and leak private user queries based on this in Section 4.
+Malicious User.
+Here we assume the attacker is a malicious user interacting with the LLM, similar to Debenedetti et al.
 (2024).
-The goal of the attacker is to leak private user queries.
-Although network packets containing LLM response tokens are encrypted, the packet size is still observable.
-The attacker can also observe the time between each packet to identify which iteration it belongs to.
-We show this allows an adversary to fingerprint and leak private user queries in Section 4.
-  2. 2.
-    Malicious User.
-In this setting, we assume the attacker is a malicious user interacting with the LLM, similar to Debenedetti et al.
-(2024), with the goal of leaking confidential data used by LLM’s speculative decoding mechanism.
-Here, the user can craft arbitrary inputs and inspect the output, and also monitor the number of tokens generated per iteration for the output tokens.
-We show this allows an adversary to leak out any confidential data used for the speculative decoding or its hyperparameters in Section 5.
+The user can craft arbitrary inputs and inspect the output, and the tokens per iteration with the goal of leaking confidential data used by the LLM’s speculative decoding mechanism for generating predictions.
+We demonstrate this attack in Section 5.
 4 Query Fingerprinting Attack
@@
 Attack Scenario.
-We consider a scenario where a user interacts with an AI chatbot using an LLM.
-The LLM is programmed to answer a predefined set of user queries through the system prompt.
-This can occur in scenarios such as a patient in a healthcare setting interacting with an LLM to get medical queries answered, or a user in a customer service setting sharing private information with the chatbot before being redirected to a human.
+Consider a medical AI chatbot, programmed to answer a predefined set of user queries through the system prompt, such as a patient in a healthcare setting asking an LLM medical symptoms related to a disease.
 The queries are sent from a client device, over a network, to a server that hosts the LLM equipped with speculative decoding.
 This leads to the threat of a network-based attacker trying to eavesdrop on the queries and responses, as described in Section 3.
-Figure 2: Overview of the Query Fingerprinting Attack.
-In the offline phase, the network-based attacker profiles the trace of number of tokens generated per iteration influenced by speculative decoding, for all possible inputs, and trains a classifier.
-In the online phase, the attacker extracts the trace of number of tokens generated per iteration for an unknown input, and uses the trained classifier to leak the input.
-Goals and Capabilities.
-Figure 2 shows the overview of our attack.
+Figure 2(a) shows the overview of our attack.
 Our attacker over the network seeks to learn the private query asked by the user.
-While the query and response packets are encrypted, the attacker can intercept the response packets, measure the time intervals between packets, and count the number of tokens per packet or simply use the packet size to approximate the number of tokens.
-Mechanism.
+While the query and response packets are encrypted, the attacker can intercept the response packets, and the packet size to approximate the number of tokens per packet.
+(a) Overview of the Query Fingerprinting Attack.
+(b) Token counts and packet sizes per iteration vs. iteration-id. for a sample response with LADE.
+Figure 2: (a) In the offline phase, the network-based attacker profiles variation in number of tokens per iteration influenced by speculative decoding, (based on packet sizes) and trains a classifier.
+In the online phase, the attacker uses the classifier to leak the input.
+(b) Packet sizes and tokens per iteration are correlated, allowing variations in packet sizes to be used to approximate token count variations.
+Attack Mechanism.
 Since the input set is limited, the attacker fingerprints inputs based on the pattern of output token counts per iteration influenced by speculative decoding.
@@
 In the online phase, the attacker collects the trace for an unknown input and uses the classifier to identify the input.
-Next, we provide more details about how these fingerprints are obtained, and then describe how the attack works on various speculative decoding techniques (LADE, REST, BiLD).
+Next, we provide more details about how these fingerprints are obtained, and then describe how the attack works on various speculative decoding techniques (LADE, REST, BiLD, and EAGLE).
 4.2 Fingerprinting Using Speculation Patterns
-Speculative decoding introduces variation in the number of tokens generated per iteration.
-In iterations with correct speculation, multiple tokens are verified in a single iteration and sent back together in a single packet.
-However, if the speculation is incorrect, a single token is generated per iteration.
-The attacker uses this variation to create a fingerprint for a particular output sequence.
-As shown in  Figure 2, from each packet containing outputs from a single iteration, the attacker measures the number of tokens generated per iteration.
-Subsequently, by plotting the token numbers vs iteration ID, it gets the trace for an input/output (prompt-response) tuple.
-Figure 3 shows the trace of tokens per iteration for two different prompts with LADE across different runs.
-From this, we observe two key properties that make these traces suitable for usage to identify and leak the prompt and response:
-  1. 1.
-    Traces are unique to a prompt and response - different prompts result in different traces of tokens per iteration, as shown in Figure 3(a) and Figure 3(b).
-  2. 2.
-    Traces are reproducible - repeating the same prompt produces similar patterns in the trace of token counts, as shown in the two iterations in Figure 3(a) or the two iterations in Figure 3(b).
-(a) What is a narcissistic personality disorder?
-(b) What are the symptoms of cancer?
-Figure 3: Trace of tokens generated per iteration vs Iteration-ID, for two runs of different prompts (a) and (b) with LADE.
-The trace has unique and reproducible patterns, enabling its usage as a fingerprint for leaking prompts.
-These observations demonstrate that the number of tokens generated per iteration with speculative decoding is a unique identifier for a prompt/response pair.
-These patterns can leak the prompt/response via side-channels to an adversary.
-However, in existing LLM deployments on servers such as OpenAI’s ChatGPT, the logits get decoded into character strings by the detokenization at the server-side before being streamed to the user in encrypted packets.
-Because the packets are encrypted, the tokens in each packet cannot be directly inferred.
-Nevertheless, there is a strong correlation between the packet size and the number of tokens per iteration.
-Figure 4 shows the trace of token count per iteration for a prompt with LADE, and the corresponding packet size in bytes per iteration match considerably.
-Thus, for our fingerprinting attack, the adversary can use the trace of packet size per iteration as a proxy for the token count per iteration, given that it is easily observable.
-Figure 4: Comparing the trace of token count per iteration and trace of packet size, versus iteration-id for a prompt response with LADE.
-Strong similarity between the traces allows the attacker to infer speculation patterns, i.e., variation in tokens per iteration, via variation in packet size.
-Next, we develop attacks to fingerprint and leak user inputs, based on the traces of packet sizes per iteration.
-4.3 Attack Experiment Design
-We perform our attacks on an AI Chatbot designed to answer common medical queries from a patient.
-The user is a patient interacting with the chatbot and the attacker seeks to learn which query the user asks.
-We conduct the following three fingerprinting attacks:
+Speculative decoding causes variation in the number of tokens generated per iteration, as correct speculation results in multiple tokens verified and sent together, while incorrect speculation generates one token per iteration.
+This variation can be used to fingerprint prompt/response pairs.
+We observe that the variations in tokens per iteration versus iteration-ID are both unique to each prompt-response combination, and are also reproducible to a great extent, as shown in Figure 4 in Appendix B.
+To quantitatively validate this, we generated 2 traces each from 50 prompts taken from MedAlpaca dataset Han et al.
+(2023) (Human-LLM interactions asking questions on different diseases; see Appendix A) input to LADE, with temperature of 0.3, and computed the pairwise cosine similarity between them.
+Traces of the same prompt (or disease) have cosine similarity of 0.9–1, whereas traces from different prompts have cosine similarity between 0.4 - 0.8.
+This shows that the trace of tokens per iteration with speculative decoding can uniquely fingerprint and identify a prompt and response.
+As packets are encrypted, the attacker cannot see the token counts directly.
+However, as shown in Figure 2(b), packet size strongly correlates with the number of tokens per iteration (Pearson Correlation Coefficient of 0.747 across all packets above).
+Thus, encrypted packet sizes serves as an effective proxy for the tokens per iteration in our fingerprinting attack.
+Note that exact token counts are not necessary; merely correlation of packet size with token counts suffices.
+Furthermore, any mitigation for prior token-size side-channels Weiss et al.
+(2024) by hiding individual token sizes is insufficient, as speculation itself introduces packet size variations due to varying token counts (see Appendix D).
+4.3 Attack Design
+We attack an AI Chatbot answering medical queries from a patient.
+The attacker seeks to learn which query a patient asks from a finite set of queries the model can answer, to potentially leak a disease the patient has, a privacy violation by law in geographies like EU or USA.
+We perform three attacks:
 Experiment 1.
-The first experiment studies a scenario where the attacker has exact knowledge of the complete set of queries, allowing the attacker to profile the exact queries offline that will be leaked out in the online phase.
-In this, we use a set of 50 prompts from real world human-LLM interactions acquired from Han et al.
+The attacker has exact knowledge of the complete set of queries, allowing the attacker to profile the exact queries offline and leak them out in the online phase.
+We choose a set of 50 prompts from a real-world human-LLM interactions dataset Han et al.
 (2023).
-These prompts ask a variety of questions about different common diseases (the list of prompts is provided in the Appendix).
+These prompts ask a variety of questions about different diseases (the list of prompts is in Appendix A).
 Experiment 2.
-In this experiment, we choose 50 prompts with commonplace diseases from our dataset that all starts with “What are the symptoms of”.
-Therefore, all the prompts have a similar structure.
-We design this experiment to see if our attack can distinguish similar prompts.
-The training and testing prompts of both Experiment 1 and 2 are the same. This illustrates the case where the attacker has exact knowledge of the set of predefined prompts.
+Next, we ask the question, do the speculation patterns depend on structural patterns or deeper semantic meanings within the response?
+To analyze this better, we choose 50 prompts about diseases that all starts with “What are the symptoms of” from the same dataset, ensuring all prompts have a similar structure.
+With this, we seek to understand whether our attack can be used to leak deeper semantic meanings associated with the prompts (e.g., diseases of the prompts) even when all the prompts have similar structures.
 Experiment 3.
-The last experiment studies a scenario where the attacker has only approximate knowledge of the set of queries asked by the user.
-To emulate this, the attacker profiles queries that are semantically similar to the queries asked by the user (e.g., profiled query: “For how long should you take zolpidem?” and user: “How long is zolpidem typically prescribed for?”).
-We reuse the prompts from Experiment 1 as our training prompts, but for testing, i.e., the user prompts to be leaked, we rephrase the prompts in the training set using ChatGPT-4o by prompting it as "Can you rephrase the following 50 questions while keeping the diseases in each question the same?"
-4.4 Methodology
+Assuming that speculation patterns rely on semantic meanings of responses, can the attacker use this to leak private information about the prompt (e.g., user’s disease) without the exact knowledge of the user’s query?
+While Experiment 1 and 2 assumed training and testing prompts to be the identical, here in Experiment 3, the attacker only has approximate knowledge of user queries during profiling.
+The attacker’s training set consists of queries that are semantically similar to user’s test-set queries (e.g., profiled query: “For how long should you take zolpidem?” and user’s query: “How long is zolpidem typically prescribed for?”).
+We use the prompts from Experiment 1 as our training prompts, but during testing, we rephrase the prompts using ChatGPT-4o by prompting it as ”Can you rephrase the following 50 questions while keeping the content in the question the same?”.
+We provide a more general experiment with an out-of-distribution training set in Section 4.8.
+4.4 Experimental Setup
 Target Models.
-We target three recent speculative decoding techniques: LADE Fu et al.
-(2024) that uses self-drafting for speculation, REST He et al.
-(2024) that uses retrieval from datastore for speculation, and BiLD Kim et al.
-(2023) which uses a smaller model for speculation.
-However, our results are as applicable to other techniques using speculation.
-We demonstrate results for LADE with TinyLlama 1.1B Chat V1.0 Zhang et al.
-(2024a), for REST with Vicuna 7B Chiang et al.
-(2023) as the target model and ShareGPT Aeala (2023) data-store containing 120K conversations used for speculation, and for BiLD with Llama2 7B Touvron et al.
-(2023) as the target model and TinyLlama 1.1B Zhang et al.
-(2024a) as the smaller model for speculation.
-We run LADE and REST on NVIDIA RTX4090 (24GB) and BiLD on NVIDIA A100 (40GB) GPUs with Question-Answer tasks, to simulate a chatbot under attack.
+We target four representative speculative decoding techniques: LADE Fu et al.
+(2024) and EAGLE Li et al.
+(2024a) using self-drafting, REST He et al.
+(2024) using retrieval and tree-verification, and BiLD Kim et al.
+(2023) using a smaller draft model.
+We run LADE with TinyLlama 1.1B Chat V1.0 Zhang et al.
+(2024a), REST with Vicuna 7B Chiang et al.
+(2023) as the target model and ShareGPT Aeala (2023) data-store of 120K conversations, and BiLD with Llama2 7B Touvron et al.
+(2023) and TinyLlama 1.1B Zhang et al.
+(2024a) as the target and draft model respectively.
+We run LADE and REST on NVIDIA RTX4090 (24GB) and BiLD on NVIDIA A100 (40GB) GPUs with Question-Answer tasks, to model a chatbot under attack.
 Datasets and Classifier.
-For each experiment, we use 50 queries of varying complexity, length, and semantics (Appendix lists all the prompts).
-For our training set, in the offline profiling phase of the attack, we run each query 5 to 30 times and extract the packet size traces (5 to 30 traces per query) and train a Random Forest classifier using them.
-For the test set in the online phase, we use the corresponding 50 queries and 5 new traces per query, collecting 250 new data points for evaluation, and report accuracy and F1-scores.
-For our attack, we use scikit-learn’s Random Forest with 150 estimators, max depth of 15, minimum 10 samples per split, minimum 1 sample per leaf, and mean squared error as the loss function.
-Accuracy (%) F1 Score
-TPQ $\rightarrow$ 5 10 20 30 5 10 20 30
+For each experiment, we use 50 queries of varying complexity, length, and semantics (Appendix A lists the prompts).
+For the training set in the profiling phase of the attack, we run each query 5 to 30 times (with varying temperatures) and extract the packet size trace (5 to 30 traces per query) and train a Random Forest classifier.
+For the test set in the online phase, we use the corresponding 50 queries and 5 new traces per query, collecting 250 data points for evaluation, and report accuracy.
+For our attack, we use scikit-learn’s Random Forest with 150 estimators, max depth of 15, minimum 10 samples per split and 1 sample per leaf, and mean squared error as loss function.
+Traces Per Query (TPQ)
+5 10 20 30
 Experiment 1
-LADE 53.6% 63.2% 65.5% 68% 0.49 0.59 0.63 0.66
-REST 91.2% 97.6% 100% 99.6% 0.91 0.98 1.0 1.0
-BiLD 28.4% 41.6% 49.2% 52% 0.27 0.43 0.52 0.55
+LADE 53.6% 63.2% 65.5% 68%
+REST 91.2% 97.6% 100% 99.6%
+BiLD 28.4% 41.6% 49.2% 52%
 Experiment 2
-LADE 33.2% 53.2% 64.8% 70.8% 0.3 0.51 0.63 0.7
-REST 96.4% 97.6% 100% 99.6% 0.96 0.98 1.0 1.0
-BiLD 58.8% 68% 75.2% 77.6% 0.54 0.64 0.73 0.76
+LADE 33.2% 53.2% 64.8% 70.8%
+REST 96.4% 97.6% 100% 99.6%
+BiLD 58.8% 68% 75.2% 77.6%
 Experiment 3
-LADE 15.2% 16% 13.6% 18.8% 0.13 0.13 0.1 0.16
-REST 35.2% 38.4% 40% 39.6% 0.27 0.32 0.34 0.33
-BiLD 13.2 % 20% 20.4% 20% 0.11 0.16 0.16 0.16
-Table 1: Results of the Query Fingerprinting Attack on three speculative decoding techniques (LADE, REST, BiLD) with temperature of 0.8, with 5 to 30 Traces Per Query (TPQ) used for training the classifier.
-4.5 Results
-Table 1 shows the results for the Query Fingerprinting Attack on LADE, REST, and BiLD.
-Exact Knowledge Scenarios (EK).
-In both Experiment 1 and 2, the attacker leaks the input prompt with high accuracy: with accuracy of up to 70.8%, 100%, and 77.6% for LADE, REST, and BiLD respectively, compared to random guessing (2% for a set of 50 prompts).
-In Experiment 1, where we study randomly selected prompts and the set of profiled and leaked prompts are identical, the attack success rates improve as the number of traces per query (TPQ) profiled increases, reaching up to 68% accuracy and an F1 score of 0.66 with 30 TPQ for LADE, reaching 99.6% accuracy with 30 TPQ and an F1 score of 1.0 for REST, and reaching 52% accuracy with 30 TPQ and an F1 score of 0.55 for BiLD.
-In Experiment 2, where all our prompts have a similar structure, the results are similar, with attack accuracy for LADE, REST, and BiLD reaching 70.8%, 99.6%, and 77.6% respectively.
-This indicates that our attack success rate is not influenced by the structure of the prompt (input), as long as the response (output) is different.
-Approximate Knowledge Scenarios (AK).
-In Experiment 3, where the leaked prompts are not exactly the same as the profiled prompts, the attack success rates are lower (up to 18.8% for LADE, 20% BiLD, and 40% for REST), but still significantly better than random-guessing (2%).
-This demonstrates that models using speculative decoding are vulnerable to query fingerprinting attacks, even when the attacker just has approximate knowledge about the queries.
-Overall, all of the attacks are more successful on REST than BiLD and LADE.
-This is because REST demonstrates higher correct speculation rates compared to BiLD and LADE, and therefore has more prompt-dependent variations in the tokens per iteration and the packet sizes, leading to the attack achieving higher accuracy on REST.
-4.6 Ablation Studies
-Table  2 presents an ablation study studying the attack accuracy as the temperature of our target model varies.
+LADE 15.2% 16% 13.6% 18.8%
+REST 35.2% 38.4% 40% 39.6%
+BiLD 13.2% 20% 20.4% 20%
+Table 1: Accuracy of the Query Fingerprinting Attack on LADE, REST, and BiLD with temperature of 0.8, using 5 to 30 Traces Per Query (TPQ) for training.
 Temperature
@@
 BiLD 17.2% 15.6% 20% 16.4%
-Table 2: Accuracy of the Query Fingerprinting Attack on LADE and REST at different temperatures (0.3, 0.6, 0.8, 1.0), with 30 Traces Per Query (TPQ) used to train the classifier.
-For LADE and BiLD, as the temperature increases from 0.3 to 1.0, the attack accuracy for both Experiment 1 and Experiment 2 decreases.
+Table 2: Ablation Study. Attack accuracy of query fingerprinting as temperatures vary (0.3, 0.6, 0.8, 1.0), using 30 Traces Per Query (TPQ) for training.
+4.5 Results
+Table 2 shows the results for the Query Fingerprinting Attack on LADE, REST, and BiLD.
+Exact Knowledge Scenarios.
+In both Experiment 1 and 2, the attacker leaks the input prompt with high accuracy: with accuracy of up to 70.8%, 100%, and 77.6% for LADE, REST, and BiLD respectively, compared to random guessing (2% for a set of 50 prompts).
+In Experiment 1, where we study randomly selected prompts and the set of profiled and leaked prompts are identical, the attack success rates improve as the number of traces per query (TPQ) profiled increases, reaching up to 68% accuracy and an F1 score of 0.66 with 30 TPQ for LADE, reaching 99.6% accuracy with 30 TPQ and an F1 score of 1.0 for REST, and reaching 52% accuracy with 30 TPQ and an F1 score of 0.55 for BiLD.
+In Experiment 2, where all our prompts have a similar structure but different semantic meaning, the results are similar, with attack accuracy for LADE, REST, and BiLD reaching 70.8%, 99.6%, and 77.6% respectively.
+This indicates that our attack success rate is not influenced by the structure of the prompt, but relies and leaks the semantic meanings of the prompt and response.
+Approximate Knowledge Scenario.
+Based on the above results, Experiment 3 uses profiled prompts that are not identical to the leaked prompts, but still semantically similar.
+Here, the attack success rates are lower (up to 18.8% for LADE, 20% BiLD, and 40% for REST), but still significantly better than random-guessing (2%).
+Thus, speculative decoding can still leak the topic of a user’s prompt (e.g., user’s disease), with queries that are approximately similar in semantics to the user queries.
+Section 4.8 details a broader out-of-distribution experiment, where the training set is taken from public sources independent of the test set.
+The attacks are more successful on REST than BiLD and LADE.
+This is because REST has higher correct speculation rates and more stable speculation patterns (due to its reliance on retrieval from a datastore of previous conversations) compared to BiLD and LADE, enabling higher attack accuracies.
+4.6 Ablation Studies
+Table  2 presents an ablation study for the attack accuracy as the temperature of our target model varies.
+For LADE and BiLD, as the temperature increases from 0.3 to 1.0, the attack accuracy for both Experiment 1 and 2 decreases.
 This is because at lower temperatures, the generations and speculation patterns are more stable and alike.
-Therefore, the output tokens for the same prompt and the packet size traces are more similar across different runs, making the attack more successful.
-For REST, the attack accuracy is relatively stable for both Experiment 1 and 2 despite the increasing temperature.
-This is likely because the features are so pronounced that it withstands the perturbations caused in the output by the increasing temperature.
-The attack accuracy in Experiment 3 has an interesting trend with temperature.
-Since the attacker only has approximate knowledge of the prompts (train and test set are different), the accuracy of this attack is much lower than Experiment 1 and 2.
-However, unlike Experiment 1 and 2, for Experiment 3 the attack accuracy has a moderate increase as the temperature of the model increases.
-This is because, at low temperatures, with limited variation in the training data, the classifier overfits on the training set, and at higher temperatures, this problem does not exist, allowing the attack to reach slightly higher accuracies.
-5 Intellectual Property Leakage
-Speculative decoding techniques may use carefully selected data or tuned hyper-parameters for achieving the highest correct speculation rates.
-In this section, we focus on how a malicious user can craft inputs to recover such data or hyper-parameters.
-We assume the malicious user can examine the output and count the number of tokens streamed back per iteration, as outlined in the threat model in Section 3-2.
-5.1 Leaking Datastore Used for Speculation
-Techniques like REST He et al.
-(2024) rely on retrieval from an external datastore to generate speculative tokens.
-These datastores can be populated with proprietary information to tailor model outputs to specific domains, or data collected from user interactions that is private.
-Patterns of correct speculations arising from such data can be exploited to leak the sensitive information in such datastores.
-5.1.1 Background on REST’s Speculation
-REST builds a datastore of phrases in the format of context and continuation pairs, where a preceding context and its following continuation are stored.
-During generation, REST retrieves the longest context matching with the tokens generated so far, and uses it to construct a speculative draft by selecting top prefixes from the retrieved continuations.
-This speculative draft is verified using a tree attention mechanism by the target model.
-5.1.2 Attack Design
-We observe that any tokens that are correctly speculated, i.e., returned as a group of tokens per iteration have to exist in the datastore.
-We propose an attack where a malicious user crafts inputs and observes which tokens are correctly speculated and all such tokens can thus be leaked from the datastore.
-With this, the attacker can systematically extract sequences from the datastore.
+Hence, the output tokens for the same prompt and the packet sizes are more similar across runs, making the attack more successful.
+For REST, the attack accuracy is stable for both Experiment 1 and 2 despite the increasing temperature.
+This is because the features are so pronounced that it withstands perturbations in the output caused by the increasing temperature.
+In Experiment 3, the attacker accuracy is lower than Exp 1 or 2, as it only has approximate knowledge of the prompts (train and test set are different) However, unlike Experiment 1 and 2, for Experiment 3 the attack accuracy increases moderately as the temperature increases.
+This is because, at low temperatures, with limited variation in the training data, the classifier overfits on the training set, and at higher temperatures, this problem does not exist, allowing the attack be slightly more accurate.
+4.7 Attack on Remote vLLM Server
+To show real-world practicality, we demonstrate our attack on a remote LLM inference server on vLLM, with Llama3 8B Instruct as the target model and EAGLE speculative decoding.
+Our LLM server is hosted on a cloud-based A100 GPU, while our client is located in an academic network more than 1500 miles away.
+Our attacker snoops the encrypted network packets in between the server and the client, using Wireshark, and captures the bytes per packet.
+We also compare our attack with prior work, Carlini & Nasr (2024), that used packet inter-arrival-times as a side-channel, that varies based on inference-time optimizations.
+To mimic their setup, which focused on ChatGPT, that sends tokens generated in the same iteration in separate packets, we modify vLLM to similarly send separate packets for individual tokens generated in the same iteration.
+For our packet size side channel, we generate our trace by combining sizes of packets received within 55 ms of each other (based on observed network latencies).
+Table 3 shows that, under no additional server load, our attack achieves a maximum of 77.6% accuracy, significantly outperforming prior work which achieves a maximum accuracy of 14.4%.
+This is because inter-arrival times provide a noisy signal due to varying network contentions.
+Moreover, under spikes in server load, the inter-arrival times (time per output token) can significantly spike as observed in recent work Zhang et al.
+(2024b), causing test-time observations to vary dramatically compared to training time.
+Consequently, assuming a standard log-normal distribution of time-per-iteration to mimic high server load, the accuracy of prior timing channel attacks drops to 6.4%.
+In comparison, our attack accuracy is unaffected, since we rely on the packet size side-channel, which remains robust to variations in server loads or network congestion.
+No Server Load High Server Load
+Temperature $\rightarrow$ 0.3 0.6 0.8 1.0 0.3 0.6 0.8 1.0
+Packet Size Side Channel (ours) 77.6% 65.6% 44.8% 24.0% 77.6% 65.6% 44.8% 24.0%
+Inter-Arrival Time Side Channel 14.4% 6.0% 5.2% 4.8% 6.4% 4% 2.4% 2.8%
+(by Carlini & Nasr (2024))
+Table 3: Accuracy of query fingerprinting attack (experiment 1) on vLLM with EAGLE’s speculative decoding at different temperatures (0.3, 0.6, 0.8, 1.0), using 30 Traces Per Query (TPQ) for training.
+We evaluate the attack using our packet-size side channel, and with inter-arrival times based side channel from prior work Carlini & Nasr (2024).
+4.8 Out-of-Distribution Training
+To study the effect of distribution mismatch between profiling and attack phases, we conducted an additional experiment where the attacker trains on an out-of-distribution (OOD) set of queries.
+Shown in  A.5, we used GPT-4o to generate 50 common diseases that users typically ask AI chatbots about, prepending the phrase “what are the symptoms of” to each.
+This OOD set was used for training, while evaluation was performed on the same 50 disease-related queries from Experiment 1 (with model temperature of 0.3 to 0.8), ensuring no prior knowledge of the test dataset during training.
+Despite the lack of prior knowledge of the true query set, the attack substantially outperformed random guessing ($\sim$2%).
+LADE achieved 23% - 25% accuracy, REST 35% - 36%, and BiLD 23% - 25%, where success was defined as predicting a disease with similar symptoms (as per GPT-5 mini) compared to the ground truth, as temperature of the target model varies between 0.3 and 0.8 (see Table  7 in Appendix E for detailed results).
+These results show that although accuracy drops relative to the in-distribution setting, speculative decoding still leaks strong and learnable signals even under OOD conditions, without being much affected by temperature, underscoring the robustness of our side-channel attack.
+5 Leaking Private Data Used for Speculation
+Speculative decoding techniques may use carefully tuned data to ensure high correct speculation rates, which may contain private user data or other intellectual property.
+In this section, we show how a malicious user can leak out private data used for predictions.
+We show this on REST He et al.
+(2024), which relies on retrieval from a datastore to generate speculative tokens.
+This datastore can be populated with proprietary information to tailor model output to specific domains, or data collected from user interactions, which may be leaked by our attack.
+A similar attack can also leak speculation hyperparameters, which can be a company’s intellectual property, as shown in Appendix C.
+Attack Design.
+REST retrieves the longest matching sequence from its datastore based on the tokens generated so far and uses the subsequent continuations as speculative tokens.
+Any tokens that are correctly speculated, i.e., returned as a group of tokens per iteration, have to exist in the datastore.
+We propose an attack where a malicious user crafts inputs and observes which tokens are correctly speculated and systematically leaks sequences of tokens from the datastore.
 We study the following attacker strategies for input generation to achieve high token leakage rates:
@@
 Leveraging Common Words.
-The LLM is prompted to generate paragraphs containing the most frequently occurring words, taken from a dataset of top 10,000 English words Brants & Franz (2021).
-This is based on the insight that there is a high chance of phrases in the datastore containing these words.
+The LLM is prompted to generate paragraphs containing the most frequently occurring words, drawn from a dataset of the 10,000 most common English words Brants & Franz (2021).
+This exploits the high likelihood that phrases in the datastore contain these words.
 3.
 Reusing Leaked Sequences.
-The LLM is prompted to generate paragraphs that extend phrases leaked out so far, as shown in Figure 5.
+The LLM is iteratively prompted to generate text extending phrases already leaked out, as shown in Figure 3(a).
 This approach is based on the insight that the sequence that is already leaked out may be a part of a longer phrase in the datastore.
-Therefore, with such iterative prompting, additional tokens from the phrase can be leaked out in a focused manner, leading to more unique tokens being leaked with a similar number of queries.
-Figure 5: Leakage process where an attacker exploits leaked sequences to extract longer datastore sequences.
-Figure 6: Unique sequences leaked from REST datastore as time increases with different attack strategies, averaged over 3 runs per strategy.
-5.1.3 Results
-We test our attacks on a REST implementation using the ShareGPT datastore of 120K conversations for speculation on NVIDIA RTX 4090.
-Figure 6 shows the unique sequences leaded by each of the three attack strategies, averaged over three runs per strategy.
-Random generation leaks 20,000 unique sequences in 20 minutes, but subsequently the rate slows down leaking only 70,000 sequences in 3 hours.
-In comparison, prompting with frequently occurring words leaks 31,000 unique sequences in 20 minutes, which increases linearly to 190,000 after 3 hours.
-Lastly, prompting with leaked sequences results in the most efficient leakage, leaking about 35,000 unique sequences in 20 minutes and about 200,000 after 3 hours, highlighting the amplification potential of feedback.
-These results show that an attacker can easily leak REST’s datastore when they can observe correct speculation patterns.
-5.2 Leakage of Speculation Hyper-parameters
-Speculative decoding mechanisms rely on hyper-parameters to control speculation, directly impacting correct speculation rates and performance.
-By analyzing traces of these patterns, we can reverse-engineer the hyper-parameters used.
-We demonstrate this attack on LADE Fu et al.
-(2024).
-5.2.1 Background on LADE’s Implementation
-LADE generates speculative tokens by caching and reusing prior n-grams generated by the model.
-LADE has two main parameters: N for n-gram size, and G for guess set size.
-Its cache is structured as a key-value store, where the key is a token, and the value is a list of G candidate (N-1)-grams following the key token, that are used for speculation.
-These are replaced in the Least-Recently-Used (LRU) order.
-During execution, LADE use the last token of the previous iteration or the input to query this cache, and the associated G candidate (N-1)-grams are used as speculation candidates, and the best match with target model generation is accepted.
-5.2.2 Leaking Hyper-parameter N in LADE
-As each iteration outputs at most N-1 tokens under correct speculation, we expect the maximum number of tokens per iteration to be N-1.
-Therefore, we prompt LADE with a simple prompt, “Repeat Letter ‘A’ 60 times” that should sustain the maximum correctly speculated tokens per iteration.
-We observe the maximum number of tokens per iteration with this prompt to deduce N-1, and thus learn N.
-Figure 7 shows the number of correct speculated tokens for LADE with three different values of $N$ for this prompt.
-We see that for each value of N (4, 5, and 6), this prompt has a maximum of 3, 4, and 5 tokens per iteration (N-1), letting us learn the value of N as 4, 5, or 6 in each of these cases.
-This technique can be extended to learn any value of N used in LADE.
-Figure 7: Number of tokens per iteration with LADE for different N, with a prompt that achieves maximum correct speculation (“Repeat Letter ‘A’ 60 times”.
-The maximum token count per iteration corresponds to value of N-1.)
-5.2.3 Leaking Hyper-parameter G in LADE
-To leak G, the number of candidate predictions considered for speculation, we force LADE to incur targeted mis-speculations.
-Here, we prompt the model to repeat a sequence of phrases with the same prefix (“I run”), i.e., “I run to the grocery shop.
-I run as fast as cars.
-I run with my best friend.
-I run from the lecture hall.".
-If the number of such phrases is less than or equal to G, LADE caches all the N-grams following the key “run”, resulting in it being able to correctly speculate the token that follows “run” in each of these phrases.
-However, if the number of phrases are greater than G, then with LRU replacement, newer n-grams evict older n-grams for the key “run” from the set of candidates which has a limited capacity of G.
-This causes the token following “run” to always be mis-speculated.
-Figure 8 shows the traces of token generation time for LADE ($G=3$), as the number of phrases ($P$) of the type “I run …” increases from 2 to 4.
-When $P$ is 2 or 3 ($\leq G$), we notice all the tokens after the initial ones are correctly predicted in the generation (high number of tokens per iteration).
-But when $P$ is 4 ($>G$), every 7th token, i.e.
-the token following “run” is mis-speculated (one token per iteration).
-Thus, we leak the value of G to be 3.
-This attack can be generalized to any value of parameter G in LADE.
-Figure 8: Time Per Token vs Token-ID for LADE with Guess Set (G = 3), using a prompt that asks the LLM to repeat P phrases sharing a common prefix “I run”.
-Periodic mis-speculations occur for the token after “run”, only when $P>G$, leaking the value of $G$.
+(a) Exploiting leaked sequences for longer extraction.
+(b) Unique Sequences Leaked vs Time (avg of 3 runs).
+Figure 3: Analysis of datastore leakage from REST.
+Results.
+We test our attacks on REST using the ShareGPT datastore of 120K conversations for retrieval on NVIDIA RTX 4090.
+Figure 3(b) shows the unique sequences leaked by each attack strategy, averaged over three runs.
+Random generation initially leaks 20K sequences in 20 minutes, but slows down, leaking only 70K sequences in 3 hours.
+Leveraging common words improves this to 31K unique sequences leaked in 20 minutes and 190K after 3 hours.
+Lastly, prompting with leaked sequences is the most effective, leaking 35K unique sequences in 20 minutes and 200K in 3 hours, demonstrating how feedback amplifies leakage.
+These results show that a malicious user can reliably extract REST’s datastore via speculation patterns.
 6 Mitigation Strategies
-In this section, we discuss possible mitigation strategies for our attacks and evaluate their impact on the attack success rate as appropriate.
 6.1 Mitigating Query Fingerprinting Attacks
-For the query fingerprinting attacks, we propose the following mitigations:
+To mitigate the query fingerprinting attacks, we propose (1) padding token packets and (2) aggregating tokens over multiple iterations. We describe both below:
 1.
@@
 The padding can be added in two ways:
-  1. 1.
-    Constant Length Padding.
-Every packet can be padded to a constant size.
-This size can be selected based on the maximum number of tokens per iteration, multiplied by the maximum bytes per token.
-We pick this size as 1024 bytes, without loss of generality.
-  2. 2.
-    Variable Length Padding.
-Every packet can be padded with a random number of bytes selected from a uniform distribution of $\epsilon\sim Unif(0,D)$ where D is a pre-defined upper bound of padding and $\epsilon$ is an integer.
-Table 3: Accuracy after padding each packet to 1024 bytes (attack: 20 traces per prompt; model temperature: 0.8)
-No Padding Padding to 1024 Bytes
+[…]
+No Padding Fixed Padding Random Padding (bytes) Packet Size Overhead
+(1024 bytes) D $\rightarrow$ 6 12 24 48 D $\rightarrow$ 6 12 24 48
 Experiment 1
-LADE 65.6% 3.2%
-REST 100% 2%
-BiLD 49.2% 2%
+[…]
+[…]
+[…]
 Experiment 2
-LADE 64.8% 2.4%
-REST 100% 2%
-BiLD 75.2% 2%
+[…]
+[…]
+[…]
 Experiment 3
-LADE 13.6% 4%
-REST 40% 2%
-BiLD 20.4% 2%
-Constant Length Padding.
-Table 3 shows the accuracy of the attack after padding every response packet to a constant size of 1024 Bytes.
-This completely mitigates the side channel, reducing the attack accuracy to close to random guessing (2%).
-However, this comes at the cost of considerable communication overhead and inefficiency, due to an increase in packet size by almost 230$\times$, as majority of the packet consists of random padding bytes.
-Table 4: Attack accuracy after padding each packets with random bytes from uniform distribution, $\epsilon\sim Unif(0,D)$, where D varies from 6 to 48.
-(attack: 20 traces per prompt; model temperature: 0.8)
-Accuracy vs Max Padding Size (D) Overhead vs D
-D $\rightarrow$ 0 6 12 24 48 6 12 24 48
-Experiment 1
-LADE 65.6% 25.6% 13.6% 8.4% 3.6% 1.5$\times$ 2$\times$ 3.1$\times$ 5.1$\times$
-REST 100% 90.4% 75.6% 48% 27.2% 1.9$\times$ 2.8$\times$ 4.6$\times$ 8.2$\times$
-BiLD 49.2% 26.8% 11.6% 5.2% 4.8% 1.7$\times$ 2.3$\times$ 3.6$\times$ 6.3$\times$
-Experiment 2
-LADE 64.8% 19.2% 10.8% 5.6% 4.4% 1.6$\times$ 2.1$\times$ 3.2$\times$ 5.4$\times$
-REST 100% 95.2% 76.8% 52% 34.4% 2$\times$ 3$\times$ 4.8$\times$ 8.7$\times$
-BiLD 75.2% 37.2% 21.6% 8.4% 5.6% 1.7$\times$ 2.4$\times$ 3.7$\times$ 6.5$\times$
-Experiment 3
-LADE 13.6% 6.8% 6% 3.6% 3.2% 1.5$\times$ 2$\times$ 3$\times$ 5$\times$
-REST 40% 29.6% 20.8% 14.8% 8.4% 1.9$\times$ 2.8$\times$ 4.6$\times$ 8.1$\times$
-BiLD 20.4% 7.6% 6.8% 6% 3.6% 1.7$\times$ 2.3$\times$ 3.6$\times$ 6.2$\times$
-Variable Length Padding.
-Table 4 shows the attack accuracy after padding responses with a variable number of bytes, $\epsilon\sim Unif(0,D)$, where D varies from 6 to 48.
-As D increases to 48, the attack accuracy for LADE and BiLD drops close to random guess (2%) for all experiments.
-The attack accuracy for REST also shows a considerable drop from 100% without padding to 27% and 34% with padding of up to 48 bytes for Experiments 1 and 2 respectively.
-Similarly, for Experiment 3, attack accuracy drops from 40% to 8% with D=48.
-These reductions in accuracy come with a 5$\times$ to 8$\times$ increase in the packet size for D=48.
-Thus, variable length padding provides a better trade-off of reduction in attack accuracy with limited increase in packet size compared to constant length padding.
+[…]
+[…]
+[…]
+Table 4: Attack accuracy under constant size padding (packets padded to 1024 bytes) and variable-length padding (random padding) mitigations.
+Variable-length padding follows a uniform distribution, $\epsilon\sim Unif(0,D)$, with $D$ from 6 to 48.
+(Attack: 20 traces per prompt; model temperature: 0.8).
+[…]
 2.
-Reduce Granularity of Attacker Observation: An alternative approach is to aggregate output tokens over multiple iterations at the server and return them to the client all together.
-This can limit observability of fine-grain speculation patterns from the packet sizes for the attacker, and reduce the effectiveness of fingerprinting attacks.
-Table 5 shows the results for the server aggregating output tokens over 3 to 20 iterations.
-The accuracy of the exact knowledge attacks (Experiment 1 and 2) for LADE, REST, and BiLD drops to a minimum of 9.2%, 52%, and 11.2% respectively after aggregating the tokens over up to 20 iterations down from 65%, 100%, and 49.2% respectively without aggregation; accuracy of the approximate knowledge attacks (Experiment 3) decreases to a minimum of 5.2%, 17.6%, and 4% respectively, down from 14%, 40%, and 20% respectively.
-Thus, reducing granularity is an effective mitigation against query fingerprinting attacks.
-One downside of this approach is that the user waits longer between updates to the output as the number of iterations the tokens are aggregated over increases.
-This limits the number of iterations tokens can be aggregated over without hindering system responsiveness.
-Table 5: Accuracy of fingerprinting attacks (20 traces per prompt and temperature of 0.8) with aggregation of tokens across 3 to 20 iterations.
-Aggregation Granularity (Number of Iterations)
-1 3 5 10 20
-Experiment 1
-LADE 65.6% 53.2% 40% 24.4% 10.8%
-REST 100% 96.8% 89.2% 80.4% 52%
-BiLD 49.2% 34% 30% 19.2% 11.2%
-Experiment 2
-LADE 64.8% 45.2% 28.8% 16.4% 9.2%
-REST 100% 97.6% 92.8% 75.6% 52%
-BiLD 75.2% 57.6% 41.6% 30% 16.8%
-Experiment 3
-LADE 13.6% 13.2% 14.4% 9.2% 5.2%
-REST 40% 33.2% 28.8% 20.4% 17.6%
-BiLD 20.4% 15.6% 10% 10% 4%
-6.2 Mitigating Datastore or Hyper-parameter Leaks
-For preventing confidential data leakage from datastore or parameter leakage attacks, in addition to the two mitigations discussed previously, we can consider one other mitigation:
-3.
-Using Only Public Data for Speculative Decoding: Any data-store used for speculation (e.g., in REST) should be constructed only with public data, and any private data or personal identifiable information (e.g., usernames, address, etc.) should be scrubbed from it before it is used.
-This minimizes the risk of sensitive data leakage through any speculation patterns that may be observable despite padding or token aggregation.
-Similarly, parameters used for speculation (such as N and G in LADE) should be assumed to be public as they can be easily leaked out.
-7 Related Work
-In this section, we describe other speculative decoding techniques that may be exploitable and related LLM performance optimizations that may lead to side-channel leakage.
-7.1 LLM Speculative Decoding Techniques
-Speculative Decoding with Tree Attention.
-SpecInfer Miao et al.
-(2024) and SpecTr Sun et al.
-(2023) observe that when a token early in the speculation draft is rejected, the entire draft gets discarded, resulting in wasted computation.
-Instead, these works propose multiple drafts per iteration organized in a tree structure to provide alternatives for speculative tokens at each position, increasing the likelihood of tokens advancing to the next step and yielding greater speedup by increasing correct speculation rates.
-These drafts are organized in a tree structure with a custom tree attention mask to prevent inter-tree interactions and verified in parallel.
-Such works (similar to REST He et al.
-(2024) which also uses tree attention) increase the correct speculation rate of tokens and introduce more input-dependent speculation patterns in the output.
-Hence these are likely to be more vulnerable to our query fingerprinting attacks, similar to REST, which is highly susceptible to our attack, as we show in this paper.
-Speculative Decoding using Target Models.
-Other works reuse the target model itself for speculative decoding by utilizing the feature layer (second-to-top layer).
-Medusa Cai et al.
-(2024) trains multiple extra heads, each responsible for a token position in the draft sequence, to generate speculative tokens.
-EAGLE and EAGLE-2 Li et al.
-(2024a; b) propose a method to perform auto-regressive decoding on the feature layers itself to generate speculative tokens.
-These techniques are conceptually no different from using an alternative smaller model for speculation like BiLD.
-Hence, they are equally susceptible to query fingerprinting attacks.
-7.2 Performance Optimizations in LLMs
-Other performance optimization techniques in LLMs that can introduce timing variation, may also be susceptible to leakage via timing side-channels.
-Adaptive KV-Cache Management.
-The KV-cache is the bottleneck for token generation latency for the memory-bound LLM inference Pope et al.
-(2023).
-Recent works, like H2O Zhang et al.
-(2024b) that performs input-dependent cache eviction or TriForceSun et al.
-(2024) that performs retrieval-based KV-cache selection for drafting may have input-dependent timing variations.
-These can introduce timing side-channels allowing query fingerprinting, similar to our attacks.
-However, as such attacks rely on timing, they are likely to be more susceptible to noise.
-Prompt Caching A significant portion of prompts sent to LLMs are duplicates, including system prompts, repeated inputs, and others.
-Prompt caching, adopted by major LLM providers such as Anthropic and OpenAI Anthropic (2024); OpenAI (2024b); Gim et al.
-(2024), precomputes attention states for these repetitive prompt components to enable reuse, and substantially reducing latency for the first token when there is reuse.
-Such techniques can leak the nature of the query based on the timing.
-Other works cache the responses for prompts with similar semantic meanings Bang (2023).
-Recent attacks Song et al.
-(2024) demonstrate query fingerprinting attacks similar to ours, when such prompt caches are shared between users.
-In contrast, our attacks are effective even without any explicit sharing of caches.
+Reduce Granularity of Attacker Observation: Alternatively, the LLM server can aggregate output tokens over multiple iterations and return them together to the client.
+This limits observability of fine-grain speculation patterns from packet sizes, reducing the success of fingerprinting attacks.
+Table 5 shows the impact of aggregating output tokens over 3 to 20 iterations.
+The accuracy of the exact knowledge attacks (Experiment 1 and 2) for LADE, REST, and BiLD drops to a minimum of 9.2%, 52%, and 11.2% respectively after aggregating tokens over 20 iterations down from 65%, 100%, and 49.2% respectively without aggregation; accuracy of the approximate knowledge attacks (Experiment 3) decreases to a minimum of 5.2%, 17.6%, and 4% respectively, down from 14%, 40%, and 20%.
+Thus, token aggregation is an effective mitigation.
+A downside of aggregating tokens over larger iterations is the increased latency between outputs, which can hinder system responsiveness.
+At the extreme, aggregating all tokens in a single packet resembles a non-streaming LLM.
+Experiment 1 Experiment 2 Experiment 3
+AGG $\rightarrow$ 1 3 5 10 20 1 3 5 10 20 1 3 5 10 20
+LADE 65.6% 53.2% 40% 24.4% 10.8% 64.8% 45.2% 28.8% 16.4% 9.2% 13.6% 13.2% 14.4% 9.2% 5.2%
+REST 100% 96.8% 89.2% 80.4% 52% 100% 97.6% 92.8% 75.6% 52% 40% 33.2% 28.8% 20.4% 17.6%
+BiLD 49.2% 34% 30% 19.2% 11.2% 75.2% 57.6% 41.6% 30% 16.8% 20.4% 15.6% 10% 10% 4%
+Table 5: Token Aggregation Mitigation. Attack accuracy (20 traces/prompt, temperature of 0.8) as aggregation granularity (AGG), i.e., iterations tokens are aggregated over, varies from 3 to 20.
+6.2 Mitigating Datastore Leaks
+Use Public Data for Speculation.
+In addition to padding or token aggregation, to prevent confidential data leakage, a simple approach is to ensure any data-store used for speculation (like in REST) only contains public data, and any private or personal identifiable information (e.g., usernames or address) is anonymized.
+This eliminates the risk of sensitive data leakage through speculation.
 7 Conclusion
 This paper reveals significant privacy risks associated with speculative decoding in Large Language Models.
-Across multiple speculative decoding techniques, we demonstrate that attacks can exploit speculation patterns to infer user inputs with $>90$% accuracy via fingerprinting attacks and also leak data from datastores used for predictions.
-These leakages highlight the careful consideration needed when deploying speculative decoding techniques to ensure that performance does not come at the cost of reduced privacy.
+Across multiple speculative decoding techniques, we demonstrate that attacks can exploit speculation patterns to infer user inputs with >90>90% accuracy via fingerprinting attacks on local models and on models remotely hosted on vLLM inference servers, and similarly leak data from datastores used for predictions.
+These leaks highlight the careful consideration needed when deploying speculative decoding techniques to ensure that performance does not come at the cost of privacy.
 Appendix A Prompts for Fingerprinting Attacks
@@
 A.3 Training prompts for Experiment 3
-Same as Section A.1
+Same as Appendix A.1
 A.4 Testing prompts for Experiment 3
@@
 What leads to the development of fatty liver?
+A.5 OOD Training prompts for Section 4.8
+What are the symptoms of Common cold?
+What are the symptoms of Influenza (flu)?
+What are the symptoms of COVID-19?
+What are the symptoms of Strep throat?
+What are the symptoms of Pneumonia?
+What are the symptoms of Bronchitis?
+What are the symptoms of Tuberculosis (TB)?
+What are the symptoms of Urinary tract infection (UTI)?
+What are the symptoms of Sexually transmitted infections (chlamydia, gonorrhea)?
+What are the symptoms of Lyme disease?
+What are the symptoms of Type 2 diabetes?
+What are the symptoms of Hypertension (high blood pressure)?
+What are the symptoms of High cholesterol?
+What are the symptoms of Obesity?
+What are the symptoms of Chronic kidney disease?
+What are the symptoms of Osteoarthritis?
+What are the symptoms of Rheumatoid arthritis?
+What are the symptoms of Asthma?
+What are the symptoms of Chronic obstructive pulmonary disease (COPD)?
+What are the symptoms of Hypothyroidism?
+What are the symptoms of Coronary artery disease?
+What are the symptoms of Heart failure?
+What are the symptoms of Atrial fibrillation?
+What are the symptoms of Stroke / transient ischemic attack (TIA)?
+What are the symptoms of Peripheral artery disease?
+What are the symptoms of Migraine?
+What are the symptoms of Epilepsy?
+What are the symptoms of Parkinson’s disease?
+What are the symptoms of Multiple sclerosis?
+What are the symptoms of Alzheimer’s disease / dementia?
+What are the symptoms of Gastroesophageal reflux disease (GERD)?
+What are the symptoms of Irritable bowel syndrome (IBS)?
+What are the symptoms of Peptic ulcer disease?
+What are the symptoms of Gallstones?
+What are the symptoms of Hepatitis (A, B, C)?
+What are the symptoms of Breast cancer?
+What are the symptoms of Prostate cancer?
+What are the symptoms of Lung cancer?
+What are the symptoms of Colorectal cancer?
+What are the symptoms of Skin cancer (melanoma, basal cell carcinoma)?
+What are the symptoms of Depression?
+What are the symptoms of Generalized anxiety disorder?
+What are the symptoms of Panic disorder?
+What are the symptoms of Bipolar disorder?
+What are the symptoms of Post-traumatic stress disorder (PTSD)?
+What are the symptoms of Eczema (atopic dermatitis)?
+What are the symptoms of Psoriasis?
+What are the symptoms of Acne?
+What are the symptoms of Rosacea?
+What are the symptoms of Fungal skin infections (ringworm, athlete’s foot)?
+Appendix B Uniqueness and Reproducibility of Fingerprints
+Figure 4 shows the trace of tokens per iteration for two different prompts with LADE (temperature of 0.3) across different runs.
+By inspecting these t races, we observe two key properties that make these traces suitable for usage to identify and leak the prompt and response:
+  1. 1.
+    Traces are unique to a prompt and response - different prompts result in different traces of tokens per iteration.
+  2. 2.
+    Traces are reproducible - repeating the same prompt produces similar patterns in the trace of token counts.
+(a) What is narcissistic personality disorder?
+(b) What are the symptoms of cancer?
+Figure 4: Trace of Tokens per iteration vs Iteration-ID, for two runs of different prompts (a) and (b) with LADE.
+The trace has unique and reproducible patterns, enabling its usage as a fingerprint for leaking prompts.
+Appendix C Leakage of Speculation Hyper-parameters
+Speculative decoding mechanisms rely on hyper-parameters to control speculation, directly impacting correct speculation rates and performance.
+By analyzing traces of these patterns, we can reverse-engineer the hyper-parameters used.
+We demonstrate this attack on LADE Fu et al.
+(2024).
+C.1 Background on LADE’s Implementation
+LADE generates speculative tokens by caching and reusing prior n-grams generated by the model.
+LADE has two main parameters: N for n-gram size, and G for guess set size.
+Its cache is structured as a key-value store, where the key is a token, and the value is a list of G candidate (N-1)-grams following the key token, that are used for speculation.
+These are replaced in the Least-Recently-Used (LRU) order.
+During execution, LADE use the last token of the previous iteration or the input to query this cache, and the associated G candidate (N-1)-grams are used as speculation candidates, and the best match with target model generation is accepted.
+C.2 Leaking Hyper-parameter N in LADE
+As each iteration outputs at most N-1 tokens under correct speculation, we expect the maximum number of tokens per iteration to be N-1.
+Therefore, we prompt LADE with a simple prompt, “Repeat Letter ‘A’ 60 times” that should sustain the maximum correctly speculated tokens per iteration.
+We observe the maximum number of tokens per iteration with this prompt to deduce N-1, and learn N.
+Figure 5(a) shows the number of correct speculated tokens for LADE with three different values of $N$ for this prompt.
+We see that for each value of N (4, 5, and 6), this prompt has a maximum of 3, 4, and 5 tokens per iteration (N-1), letting us learn the value of N as 4, 5, or 6 in each of these cases.
+This technique can be extended to learn any value of N used in LADE.
+(a) Tokens per iteration with LADE for different NN, with a prompt that achieves maximum correct speculation (“Repeat Letter ‘A’ 60 times”). The maximum tokens per iteration corresponds to N-1.
+(b) Tokens per iteration with LADE with Guess Set (G = 3), using a prompt that asks to repeat P phrases sharing a common prefix “I run”.
+Periodic mis-speculations occur after “run”, only when $P>G$, leaking the value of $G$.
+Figure 5: Leaking Speculation Hyper-parameters of LADE.
+C.3 Leaking Hyper-parameter G in LADE
+To leak G, the number of candidate predictions considered for speculation, we force LADE to incur targeted mis-speculations.
+Here, we prompt the model to repeat a sequence of phrases with the same prefix (“I run”), i.e., “I run to the grocery shop.
+I run as fast as cars.
+I run with my best friend.
+I run from the lecture hall.”.
+If the number of such phrases is less than or equal to G, LADE caches all the N-grams following the key “run”, resulting in it being able to correctly speculate the token that follows “run” in each of these phrases.
+However, if the number of phrases are greater than G, then with LRU replacement, newer n-grams evict older n-grams for the key “run” from the set of candidates which has a limited capacity of G.
+This causes the token following “run” to always be mis-speculated.
+Figure 5(b) shows the traces of token generation time for LADE ($G=3$), as the number of phrases ($P$) of the type “I run …” increases from 2 to 4.
+When $P$ is 2 or 3 ($\leq G$), we notice all the tokens after the initial ones are correctly predicted in the generation (high number of tokens per iteration).
+But when $P$ is 4 (>G>G), every 7th token, i.e.
+the token following “run” is mis-speculated (one token per iteration).
+Thus, we leak the value of G to be 3.
+This attack can be generalized to any value of parameter G in LADE.
+Appendix D Impact of Mitigation for Token-Length Side Channel
+In this experiment, we want to measure the impact on our attacks of mitigation against prior attacks, such as token-length side channels Weiss et al.
+(2024) that leaks the value of a token based on its associated character length.
+We evaluate the impact of a mitigation, where every token has been padded to the same size to mask its character length.
+We measure the attack success rate for our attacks with this mitigation.
+Note that even with this mitigation, there is variation in the packet sizes, due to simply the number of tokens contained in each packet, which varies due to speculation.
+Thus, measuring the packet sizes in this scenario is the same as measuring the token counts.
+[…]
+Temperature
+0.3 0.6 0.8 1.0
+Experiment 1
+LADE 51.6% 38% 27.6% 22.8%
+REST 24% 17.2% 10.8% 8.4%
+BiLD 67.6% 43.2% 30.4% 22.8%
+Experiment 2
+LADE 41.2% 24.8% 22.4% 11.2%
+REST 25.6% 14% 14.4% 10%
+BiLD 82.4% 62.4% 46.8% 33.2%
+Experiment 3
+LADE 8% 6.8% 6.8% 3.2%
+REST 6.4% 4.8% 8% 5.2%
+BiLD 8.4% 10% 10.8% 6.8%
+Table 6: Accuracy of the Query Fingerprinting Attack with each token padded to the maximum token size (1024 bytes)
+Appendix E Out of Distribution Training
+Temperature $\rightarrow$ 0.3 0.6 0.8
+LADE 25% 24% 23%
+REST 36% 35% 36%
+BiLD 25% 23% 24%
+Table 7: Accuracy of the Query Fingerprinting Attack with out of distribution training, as the temperature of target model varies from 0.3 to 0.8.

### v4

@@
 In this work, we reveal a new side-channel whereby input-dependent patterns of correct and incorrect speculations can be inferred by monitoring per-iteration token counts or packet sizes.
-We demonstrate that an adversary observing these patterns can fingerprint user queries with ¿90% accuracy across four speculative-decoding schemes, REST ($\sim$100%), LADE (up to 92%), BiLD (up to 95%), and EAGLE (up to 77.6%) and leak confidential datastore contents used for prediction at rates exceeding 25 tokens/sec.
-We evaluate the side-channel attacks in both research prototypes as well as the production-grade vLLM serving framework.
+In evaluations using research prototypes and production-grade vLLM serving frameworks, we show that an adversary monitoring these patterns can fingerprint user queries (from a set of 50 prompts) with over 75% accuracy across four speculative-decoding schemes at temperature 0.3: REST (100%), LADE (91.6%), BiLD (95.2%), and EAGLE (77.6%).
+Even at temperature 1.0, accuracy remains far above the 2% random baseline—REST (99.6%), LADE (61.2%), BiLD (63.6%), and EAGLE (24%).
+We also show the capability of the attacker to leak confidential datastore contents used for prediction at rates exceeding 25 tokens/sec.
 To defend against these, we propose and evaluate a suite of mitigations, including packet padding and iteration-wise token aggregation.
 1 Introduction
-Large Language Models (LLMs) have transformed natural language processing (NLP), allowing machines to generate and understand human language at an unprecedented scale Vaswani et al.
-(2017); Devlin et al.
-(2019); Brown et al.
-(2020); Zhang et al.
-(2022); Le Scao et al.
-(2022); Touvron et al.
-(2023).
-LLMs typically generate text using auto-regressive decoding Touvron et al.
-(2023); OpenAI (2024), where the generation happens serially and each token depends on all the previous ones.
-Unfortunately, this serial process causes a significant bottleneck in the LLM response latency Miao et al.
-(2024); Fu et al.
-(2024) and under-utilizes the available hardware-level parallelism, limiting token generation throughput and latency.
+Large Language Models (LLMs) have transformed natural language processing (NLP), allowing machines to generate and understand human language at an unprecedented scale (Vaswani et al., 2017; Devlin et al., 2019; Brown et al., 2020; Zhang et al., 2022; Le Scao et al., 2022; Touvron et al., 2023).
+LLMs typically generate text using auto-regressive decoding (Touvron et al., 2023; OpenAI, 2024), where the generation happens serially and each token depends on all the previous ones.
+Unfortunately, this serial process causes a significant bottleneck in the LLM response latency (Miao et al., 2024; Fu et al., 2024) and under-utilizes the available hardware-level parallelism, limiting token generation throughput and latency.
 […]
 However, the adoption of speculation techniques is not without risks.
-Speculative execution in CPUs  Burton (1985); Hennessy & Patterson (2012), which inspired speculative decoding, has led to security vulnerabilities in processors, such as Spectre Kocher et al.
-(2019) and their variants, which exploit side-channels, i.e.
+Speculative execution in CPUs  (Burton, 1985; Hennessy & Patterson, 2012), which inspired speculative decoding, has led to security vulnerabilities in processors, such as Spectre (Kocher et al., 2019) and their variants, which exploit side-channels, i.e.
 timing variations due to mis-speculations that leak secret data accessed during mis-speculations.
@@
 (c) Using these patterns, a network-based adversary can fingerprint user queries and learn private user prompts and responses; a malicious user can observe correct predictions and leak out data-stores and hyper-parameters used for predictions.
-Figure 1(a) shows a response by TinyLlama 1.1B Chat  Zhang et al.
-(2024a) with a speculative decoding technique, Lookahead Decoding (LADE) Fu et al.
-(2024), with correct (blue) and mis-speculated (red) tokens illustrated.
+Figure 1(a) shows a response by TinyLlama 1.1B Chat  (Zhang et al., 2024a) with a speculative decoding technique, Lookahead Decoding (LADE) (Fu et al., 2024), with correct (blue) and mis-speculated (red) tokens illustrated.
 On correct speculation, up to $n$ tokens are verified per iteration, as shown in Figure 1(b), while on mis-speculation, only one token is generated per iteration.
-In LLM applications streaming responses to the user, token by token, a network-based adversary, such as a malicious ISP or a compromised router Weiss et al.
-(2024), can observe network packet sizes corresponding to each iteration to deduce the number of tokens per iteration and learn the input-dependent pattern of correct and incorrect speculations.
-We target medical chatbots (e.g., Hippocratic AI Shah (2025)), where users interact by asking questions about symptoms or diseases, as this is a privacy-sensitive application where any information leakage (e.g., disease or symptoms of users) breaks privacy laws.
+In LLM applications streaming responses to the user, token by token, a network-based adversary, such as a malicious ISP or a compromised router (Weiss et al., 2024), can observe network packet sizes corresponding to each iteration to deduce the number of tokens per iteration and learn the input-dependent pattern of correct and incorrect speculations.
+We target medical chatbots (e.g., Hippocratic AI (Shah, 2025)), where users interact by asking questions about symptoms or diseases, as this is a privacy-sensitive application where any information leakage (e.g., disease or symptoms of users) breaks privacy laws.
 Exploits.
-We demonstrate privacy breaches based on speculation patterns, as shown in Figure 1(c).
+We demonstrate privacy breaches based on speculation patterns (Figure 1(c)).
 First, we show a query fingerprinting attack (e.g., leaking diseases or symptoms in medical chatbots) by a network adversary.
-By profiling speculation pattern fingerprints for prompt-response pairs offline, an adversary can then use compare the speculation pattern for an unknown query with the fingerprints to recover private user prompts.
-Across speculative decoding schemes (BiLD, REST, LADE) Fu et al.
-(2024); He et al.
-(2024); Kim et al.
-(2023), our prompt identification attacks reaches accuracies of $\sim$100% for REST, 92% for LADE, and 95% for BiLD with only 30 training samples.
-On a remote vLLM inference server Kwon et al.
-(2023), which supports speculation with EAGLE Li et al.
-(2024a), we similarly observe a high accuracy of 77.6%.
+By profiling speculation pattern fingerprints for prompt-response pairs offline, an adversary can then compare the speculation pattern for an unknown query with the fingerprints to recover private user prompts.
+Across speculative decoding schemes (BiLD, REST, LADE) (Fu et al., 2024; He et al., 2024; Kim et al., 2023), our prompt identification attacks (within a set of 50 queries) reaches accuracies of $\sim$100% for REST, 91.6% for LADE, and 95% for BiLD with a temperature of 0.3.
+On a remote vLLM inference server (Kwon et al., 2023), which supports speculation with EAGLE (Li et al., 2024a), we similarly observe a high accuracy of 77.6%.
 We further show that when the actual prompts are unavailable, the adversary can train effective fingerprints using publicly available proxy datasets; e.g., by querying a medical chatbot with the top 50 most common diseases as stand-ins for real prompts.
-In this case, the attack still achieves 20–40% accuracy in recovering the disease or symptoms from user prompts, far exceeding random guessing (2%).
-Lastly, we show how adversaries crafting malicious inputs can also leak confidential data such as data-store contents used for predictions with REST He et al.
-(2024) at a rate of more than 25 tokens/second by observing correctly predicted tokens.
+In this case, the attack still achieves 20–40% accuracy in recovering the disease or symptoms from user prompts, far exceeding random guessing (2-6%).
+Lastly, we show how adversaries crafting malicious inputs can leak confidential data-store contents used for predictions with REST (He et al., 2024) at a rate of >>25 tokens/second by observing correctly predicted tokens.
 […]
@@
 Speculation with Smaller Draft Models.
-Prior works propose using smaller, faster draft models to generate sequences of tokens speculatively Miao et al.
-(2024); Kim et al.
-(2023), and then verify a single sequence or a tree of such sequences in parallel using the larger target model.
-The draft model can be a smaller version of the model family or a pruned larger model Yan et al.
-(2025).
-In this paper, we demonstrate our attacks on BiLD Kim et al.
-(2023), as a example of this type of speculation.
+Prior works propose using smaller, faster draft models to generate sequences of tokens speculatively (Miao et al., 2024; Kim et al., 2023), and then verify a single sequence or a tree of such sequences in parallel using the larger target model.
+The draft model can be a smaller version of the model family or a pruned larger model (Yan et al., 2025).
+In this paper, we demonstrate our attacks on BiLD (Kim et al., 2023), as a example of this type of speculation.
 Speculation via Self-Drafting.
 Recent works reuse the target model for producing predictions, eliminating the need for a separate draft model.
-Look-Ahead Decoding (LADE) Fu et al.
-(2024) caches previous output tokens and uses them to speculate future tokens via a key-value store populated with token sequence N-grams.
+Look-Ahead Decoding (LADE) (Fu et al., 2024) caches previous output tokens and uses them to speculate future tokens via a key-value store populated with token sequence N-grams.
 These are populated using Jacobi decoding, by selecting random tokens from the input and generating successive chain of tokens.
 When this key token reappears in the output stream, LADE uses the stored N-grams for speculation.
-Medusa Cai et al.
-(2024) trains extra heads in the target model, each responsible for a token position in the draft sequence, for speculative tokens.
-EAGLE and EAGLE-2 Li et al.
-(2024a; b) perform auto-regressive decoding directly on the feature layers itself to generate speculative tokens, and are integrated in the vLLM inference serving platform.
+Medusa (Cai et al., 2024) trains extra heads in the target model, each responsible for a token position in the draft sequence, for speculative tokens.
+EAGLE and EAGLE-2 (Li et al., 2024a; b) perform auto-regressive decoding directly on the feature layers itself to generate speculative tokens, and are integrated in the vLLM inference serving platform.
 Our attacks target LADE (locally) and EAGLE (on a remote vLLM server) as representative examples, but our techniques are broadly applicable to other techniques as well.
 Speculation Using Retrieval.
-REST He et al.
-(2024) generates predictions by retrieving data from a datastore containing relevant text or code, organized as a set of prefixes and continuations that can be potentially used as speculative tokens.
+REST (He et al., 2024) generates predictions by retrieving data from a datastore containing relevant text or code, organized as a set of prefixes and continuations that can be potentially used as speculative tokens.
 If there are multiple candidates, REST organizes them in a tree, applying a tree attention mask during decoding with the target model.
 Tree-Based Verification.
-SpecInfer Miao et al.
-(2024), SpecTr Sun et al.
-(2023) and REST He et al.
-(2024) generate multiple drafts of speculation per iteration, organized in a tree structure, to increase the length of successful verification for higher speedup.
+SpecInfer (Miao et al., 2024), SpecTr (Sun et al., 2023) and REST (He et al., 2024) generate multiple drafts of speculation per iteration, organized in a tree structure, to increase the length of successful verification for higher speedup.
 These drafts are organized with a custom tree attention mask to prevent inter-tree interactions and verified in parallel.
-In this work, we demonstrate our attacks on REST He et al.
-(2024), as a representative example of such techniques.
+In this work, we demonstrate our attacks on REST (He et al., 2024), as a representative example of such techniques.
 2.2 Side-Channel Attacks on LLMs
-Debenedetti et al.
-(2024) pioneered side-channel attacks on LLMs.
+(Debenedetti et al., 2024) pioneered side-channel attacks on LLMs.
 Their attack exploits variation in output token counts due to the tokenizer, to leak uncommon strings in the tokenizer vocabulary.
 Our work instead exploits variation in token count per iteration due to speculative decoding.
-Weiss et al.
-(2024) identified a token-length side channel in streaming LLMs, where network attackers can infer token values by analyzing encrypted packet sizes.
+(Weiss et al., 2024) identified a token-length side channel in streaming LLMs, where network attackers can infer token values by analyzing encrypted packet sizes.
 This attack reconstructs around 27% of the LLM outputs using the character length of each token.
@@
 Moreover, we show mitigating the token-length side channel by padding tokens to the maximum token-size still results in packet size variations due to speculative decoding varying the output token counts per packet.
-We show that this can be used to fingerprint prompts and responses in our attacks with greater than 50% accuracy, even after the token-length side-channel is mitigated, in Appendix D.
-Carlini & Nasr (2024) demonstrated timing attacks on LLM inference optimizations in commercial models like GPT-4 and Claude.
+We show that this can be used to fingerprint prompts and responses in our attacks with greater than 50% accuracy, even after the token-length side-channel is mitigated, in Appendix E.
+(Carlini & Nasr, 2024) demonstrated timing attacks on LLM inference optimizations in commercial models like GPT-4 and Claude.
 Their attacks rely on timing variations due to these optimizations, i.e.
-variations in packet inter-arrival times, which is considerably susceptible to noise due to network congestion wand server loads.
+variations in packet inter-arrival times, which is considerably susceptible to noise due to network congestion and server loads.
 In comparison, our attacks exploit packet size variations due to inference optimizations, which are inherently more robust and independent of server or network load variations.
 We show the resilience of our attacks compared to this prior work in Section 4.7.
-Song et al.
-(2024) and  Zheng et al.
-(2024) demonstrated timing side-channel attacks exploiting shared key-value (KV) and semantic caches to leak sensitive LLM inputs.
-Wiretapping LLMs Soleimani et al.
-(2025) also reveals private attributes in LLM-based medical and financial applications.
+(Song et al., 2024) and  (Zheng et al., 2024) demonstrated timing side-channel attacks exploiting shared key-value (KV) and semantic caches to leak sensitive LLM inputs.
+Wiretapping LLMs (Soleimani et al., 2025) also reveals private attributes in LLM-based medical and financial applications.
 3 Threat Model
-Like prior works Weiss et al. (2024), we focus on streaming LLMs, where the LLM server sends back responses to the user token-by-token. Here, we study the following attacker models.
+Like prior works (Weiss et al., 2024), we focus on streaming LLMs, where the LLM server sends back responses to the user token-by-token. Here, we study the following attacker models.
 Network Attacker.
-We assume a malicious ISP or a compromised router that can observe network packet sizes, similar to Weiss et al.
-(2024).
+We assume a malicious ISP or a compromised router that can observe network packet sizes, similar to  (Weiss et al., 2024).
 The attacker’s goal is to leak private user queries.
@@
 Malicious User.
-Here we assume the attacker is a malicious user interacting with the LLM, similar to Debenedetti et al.
-(2024).
+Here we assume the attacker is a malicious user interacting with the LLM, similar to  (Debenedetti et al., 2024).
 The user can craft arbitrary inputs and inspect the output, and the tokens per iteration with the goal of leaking confidential data used by the LLM’s speculative decoding mechanism for generating predictions.
@@
 We observe that the variations in tokens per iteration versus iteration-ID are both unique to each prompt-response combination, and are also reproducible to a great extent, as shown in Figure 8 in Appendix B.
-To quantitatively validate this, we generated 2 traces each from 50 prompts taken from MedAlpaca dataset Han et al.
-(2023) (Human-LLM interactions asking questions on different diseases; see Appendix A) input to LADE, with temperature of 0.3, and computed the pairwise cosine similarity between them.
+To quantitatively validate this, we generated 2 traces each from 50 prompts taken from MedAlpaca dataset (Han et al., 2023) (Human-LLM interactions asking questions on different diseases; see Appendix A) input to LADE, with temperature of 0.3, and computed the pairwise cosine similarity between them.
 Traces of the same prompt (or disease) have cosine similarity of 0.9–1, whereas traces from different prompts have cosine similarity between 0.4 - 0.8.
@@
 Note that exact token counts are not necessary; merely correlation of packet size with token counts suffices.
-Furthermore, any mitigation for prior token-size side-channels Weiss et al.
-(2024) by hiding individual token sizes is insufficient, as speculation itself introduces packet size variations due to varying token counts (see Appendix D).
+Furthermore, any mitigation for prior token-size side-channels (Weiss et al., 2024) by hiding individual token sizes is insufficient, as speculation itself introduces packet size variations due to varying token counts (see Appendix E).
 4.3 Attack Design
@@
 We perform three attacks:
-Experiment 1.
-The attacker has exact knowledge of the complete set of queries, allowing the attacker to profile the exact queries offline and leak them out in the online phase.
-We choose a set of 50 prompts from a real-world human-LLM interactions dataset Han et al.
-(2023).
+Experiment 1 - Exact Knowledge The attacker has exact knowledge of the complete set of queries, allowing the attacker to profile the exact queries offline and leak them out in the online phase.
+We choose a set of 50 prompts from a real-world human-LLM interactions dataset (Han et al., 2023).
 These prompts ask a variety of questions about different diseases (the list of prompts is in Appendix A).
-Experiment 2.
-Next, we ask the question, do the speculation patterns depend on structural patterns or deeper semantic meanings within the response?
+Experiment 2 - Exact Knowledge (Constrained) Next, we ask the question, do the speculation patterns depend on structural patterns or deeper semantic meanings within the response?
 To analyze this better, we choose 50 prompts about diseases that all starts with “What are the symptoms of” from the same dataset, ensuring all prompts have a similar structure.
 With this, we seek to understand whether our attack can be used to leak deeper semantic meanings associated with the prompts (e.g., diseases of the prompts) even when all the prompts have similar structures.
-Experiment 3.
-Assuming that speculation patterns rely on semantic meanings of responses, can the attacker use this to leak private information about the prompt (e.g., user’s disease) without the exact knowledge of the user’s query?
+Experiment 3 - Approximate Knowledge Assuming that speculation patterns rely on semantic meanings of responses, can the attacker use this to leak private information about the prompt (e.g., user’s disease) without the exact knowledge of the user’s query?
 While Experiment 1 and 2 assumed training and testing prompts to be the identical, here in Experiment 3, the attacker only has approximate knowledge of user queries during profiling.
@@
 Target Models.
-We target four representative speculative decoding techniques: LADE Fu et al.
-(2024) and EAGLE Li et al.
-(2024a) using self-drafting, REST He et al.
-(2024) using retrieval and tree-verification, and BiLD Kim et al.
-(2023) using a smaller draft model.
-We run LADE with TinyLlama 1.1B Chat V1.0 Zhang et al.
-(2024a), REST with Vicuna 7B Chiang et al.
-(2023) as the target model and ShareGPT Aeala (2023) data-store of 120K conversations, and BiLD with Llama2 7B Touvron et al.
-(2023) and TinyLlama 1.1B Zhang et al.
-(2024a) as the target and draft model respectively.
+We target four representative speculative decoding techniques: LADE (Fu et al., 2024) and EAGLE (Li et al., 2024a) using self-drafting, REST (He et al., 2024) using retrieval and tree-verification, and BiLD (Kim et al., 2023) using a smaller draft model.
+We run LADE with TinyLlama 1.1B Chat V1.0 (Zhang et al., 2024a), REST with Vicuna 7B (Chiang et al., 2023) as the target model and ShareGPT (Aeala, 2023) data-store of 120K conversations, and BiLD with Llama2 7B (Touvron et al., 2023) and TinyLlama 1.1B (Zhang et al., 2024a) as the target and draft model respectively.
 We run LADE and REST on NVIDIA RTX4090 (24GB) and BiLD on NVIDIA A100 (40GB) GPUs with Question-Answer tasks, to model a chatbot under attack.
@@
 For each experiment, we use 50 queries of varying complexity, length, and semantics (Appendix A lists the prompts).
-For the training set in the profiling phase of the attack, we run each query 5 to 30 times (with varying temperatures) and extract the packet size trace (5 to 30 traces per query) and train a Random Forest classifier.
+For our attack, we studied three types of classifiers - CNN, Gaussian Mixture Models (GMM) and Random Forest, and use Random Forest since it has the best accuracy.
+(See Figure 12 in Appendix G for the accuracy comparisons of Random Forest with CNN and GMM).
+For the training set in the profiling phase of the attack, we run each query 5 to 30 times (with varying temperatures) and extract the packet size trace (5 to 30 traces per query) to train our Random Forest classifier.
 For the test set in the online phase, we use the corresponding 50 queries and 5 new traces per query, collecting 250 data points for evaluation, and report accuracy.
 For our attack, we use scikit-learn’s Random Forest with 150 estimators, max depth of 15, minimum 10 samples per split and 1 sample per leaf, and mean squared error as loss function.
-Traces Per Query (TPQ)
-5 10 20 30
-Experiment 1
-LADE 53.6% 63.2% 65.5% 68%
-REST 91.2% 97.6% 100% 99.6%
-BiLD 28.4% 41.6% 49.2% 52%
-Experiment 2
-LADE 33.2% 53.2% 64.8% 70.8%
-REST 96.4% 97.6% 100% 99.6%
-BiLD 58.8% 68% 75.2% 77.6%
-Experiment 3
-LADE 15.2% 16% 13.6% 18.8%
-REST 35.2% 38.4% 40% 39.6%
-BiLD 13.2% 20% 20.4% 20%
-Table 1: Accuracy of the Query Fingerprinting Attack on LADE, REST, and BiLD with temperature of 0.8, using 5 to 30 Traces Per Query (TPQ) for training.
-Temperature
-0.3 0.6 0.8 1.0
-Experiment 1
-LADE 91.6% 85.6% 68% 61.2%
-REST 100% 100% 99.6% 99.6%
-BiLD 74.4% 59.2% 52% 34.8%
-Experiment 2
-LADE 90.8% 84.4% 70.8% 60.4%
-REST 99.6% 100% 99.6% 100%
-BiLD 95.2% 88.4% 77.6% 63.6%
-Experiment 3
-LADE 16.8% 17.6% 18.8% 19.6%
-REST 35.2% 42% 39.6% 43.2%
-BiLD 17.2% 15.6% 20% 16.4%
-Table 2: Ablation Study. Attack accuracy of query fingerprinting as temperatures vary (0.3, 0.6, 0.8, 1.0), using 30 Traces Per Query (TPQ) for training.
+Figure 3: Accuracy of the query fingerprinting attack on LADE, REST, and BiLD with temperature of 0.8, using 5 to 30 Traces Per Query (TPQ) for training.
 4.5 Results
-Table 2 shows the results for the Query Fingerprinting Attack on LADE, REST, and BiLD.
+Figure 3 shows the results for the query fingerprinting attack on LADE, REST, and BiLD.
 Exact Knowledge Scenarios.
 In both Experiment 1 and 2, the attacker leaks the input prompt with high accuracy: with accuracy of up to 70.8%, 100%, and 77.6% for LADE, REST, and BiLD respectively, compared to random guessing (2% for a set of 50 prompts).
-In Experiment 1, where we study randomly selected prompts and the set of profiled and leaked prompts are identical, the attack success rates improve as the number of traces per query (TPQ) profiled increases, reaching up to 68% accuracy and an F1 score of 0.66 with 30 TPQ for LADE, reaching 99.6% accuracy with 30 TPQ and an F1 score of 1.0 for REST, and reaching 52% accuracy with 30 TPQ and an F1 score of 0.55 for BiLD.
-In Experiment 2, where all our prompts have a similar structure but different semantic meaning, the results are similar, with attack accuracy for LADE, REST, and BiLD reaching 70.8%, 99.6%, and 77.6% respectively.
+In Experiment 1, where the set of profiled and leaked prompts are identical, the attack success rates improve as the number of traces per query (TPQ) profiled increases, reaching up to 68% accuracy and an F1 score of 0.66 with 30 TPQ for LADE, reaching 99.6% accuracy with 30 TPQ and an F1 score of 1.0 for REST, and reaching 52% accuracy with 30 TPQ and an F1 score of 0.55 for BiLD.
+In Experiment 2, where all our prompts have a similar structure but different semantic meaning, the results are similar, with attack accuracy for LADE, REST, and BiLD reaching 90.8%, 99.6%, and 95.2% respectively.
 This indicates that our attack success rate is not influenced by the structure of the prompt, but relies and leaks the semantic meanings of the prompt and response.
@@
 4.6 Ablation Studies
-Table  2 presents an ablation study for the attack accuracy as the temperature of our target model varies.
+Figure 4 presents an ablation study for the attack accuracy as the temperature of our target model varies.
 For LADE and BiLD, as the temperature increases from 0.3 to 1.0, the attack accuracy for both Experiment 1 and 2 decreases.
@@
 This is because the features are so pronounced that it withstands perturbations in the output caused by the increasing temperature.
+Figure 4: Attack accuracy of query fingerprinting as temperatures vary (0.3, 0.6, 0.8, 1.0), using 30 Traces Per Query (TPQ) for training.
 In Experiment 3, the attacker accuracy is lower than Exp 1 or 2, as it only has approximate knowledge of the prompts (train and test set are different) However, unlike Experiment 1 and 2, for Experiment 3 the attack accuracy increases moderately as the temperature increases.
@@
 Our attacker snoops the encrypted network packets in between the server and the client, using Wireshark, and captures the bytes per packet.
-We also compare our attack with prior work, Carlini & Nasr (2024), that used packet inter-arrival-times as a side-channel, that varies based on inference-time optimizations.
+We also compare our attack with prior work, (Carlini & Nasr, 2024), that used packet inter-arrival-times as a side-channel, that varies based on inference-time optimizations.
 To mimic their setup, which focused on ChatGPT, that sends tokens generated in the same iteration in separate packets, we modify vLLM to similarly send separate packets for individual tokens generated in the same iteration.
 For our packet size side channel, we generate our trace by combining sizes of packets received within 55 ms of each other (based on observed network latencies).
-Table 3 shows that, under no additional server load, our attack achieves a maximum of 77.6% accuracy, significantly outperforming prior work which achieves a maximum accuracy of 14.4%.
+No Server Load High Server Load
+[…]
+Random Forest Packet Size (ours) 77.6% 65.6% 44.8% 24.0% 77.6% 65.6% 44.8% 24.0%
+Inter-Arrival Time 14.4% 6.0% 5.2% 4.8% 6.4% 4% 2.4% 2.8%
+(Carlini & Nasr, 2024)
+GMM Packet Size (ours) 27.2% 7.2% 4.8% 4% 27.2% 7.2% 4.8% 4%
+Inter-Arrival Time 2.6% 1.2% 3.2% 1.2% 1.6% 1.2% 2.8% 1.2%
+(Carlini & Nasr, 2024)
+Table 1: Attack accuracy for the query fingerprinting (Experiment 1 - Exact Knowledge) on vLLM with EAGLE’s speculative decoding at different temperatures (0.3, 0.6, 0.8, 1.0), using 30 Traces Per Query (TPQ) for training.
+We evaluate the attack using our packet-size side channel, and with inter-arrival times based side channel from prior work (Carlini & Nasr, 2024).
+Table 1 shows that, under no additional server load, our attack achieves a maximum of 77.6% accuracy, significantly outperforming prior work which achieves a maximum accuracy of 14.4% (both using Random Forest classifier).
 This is because inter-arrival times provide a noisy signal due to varying network contentions.
-Moreover, under spikes in server load, the inter-arrival times (time per output token) can significantly spike as observed in recent work Zhang et al.
-(2024b), causing test-time observations to vary dramatically compared to training time.
-Consequently, assuming a standard log-normal distribution of time-per-iteration to mimic high server load, the accuracy of prior timing channel attacks drops to 6.4%.
+Moreover, under spikes in server load, the inter-arrival times (time per output token) can significantly spike as observed in recent work (Zhang et al., 2024b), causing test-time observations to vary dramatically compared to training time.
+Consequently, assuming a standard log-normal distribution of time-per-iteration to mimic high server load  (Benson et al., 2009), the accuracy of prior timing channel attacks drops to 6.4%.
 In comparison, our attack accuracy is unaffected, since we rely on the packet size side-channel, which remains robust to variations in server loads or network congestion.
-No Server Load High Server Load
-Temperature $\rightarrow$ 0.3 0.6 0.8 1.0 0.3 0.6 0.8 1.0
-Packet Size Side Channel (ours) 77.6% 65.6% 44.8% 24.0% 77.6% 65.6% 44.8% 24.0%
-Inter-Arrival Time Side Channel 14.4% 6.0% 5.2% 4.8% 6.4% 4% 2.4% 2.8%
-(by Carlini & Nasr (2024))
-Table 3: Accuracy of query fingerprinting attack (experiment 1) on vLLM with EAGLE’s speculative decoding at different temperatures (0.3, 0.6, 0.8, 1.0), using 30 Traces Per Query (TPQ) for training.
-We evaluate the attack using our packet-size side channel, and with inter-arrival times based side channel from prior work Carlini & Nasr (2024).
+Even when using GMM instead of Random Forest like prior work (Carlini & Nasr, 2024), the trends are similar, with our packet-size side-channel outperforming prior inter-arrival time side-channel (Carlini & Nasr, 2024) under both low and high server loads.
 4.8 Out-of-Distribution Training
 To study the effect of distribution mismatch between profiling and attack phases, we conducted an additional experiment where the attacker trains on an out-of-distribution (OOD) set of queries.
-Shown in  A.5, we used GPT-4o to generate 50 common diseases that users typically ask AI chatbots about, prepending the phrase “what are the symptoms of” to each.
+Shown in Appendix A.5, we used GPT-4o to generate 50 common diseases that users typically ask AI chatbots about, prepending the phrase “what are the symptoms of” to each.
 This OOD set was used for training, while evaluation was performed on the same 50 disease-related queries from Experiment 1 (with model temperature of 0.3 to 0.8), ensuring no prior knowledge of the test dataset during training.
-Despite the lack of prior knowledge of the true query set, the attack substantially outperformed random guessing ($\sim$2%).
-LADE achieved 23% - 25% accuracy, REST 35% - 36%, and BiLD 23% - 25%, where success was defined as predicting a disease with similar symptoms (as per GPT-5 mini) compared to the ground truth, as temperature of the target model varies between 0.3 and 0.8 (see Table  7 in Appendix E for detailed results).
+Despite the lack of prior knowledge of the true query set, the attack substantially outperformed our previous naive random guessing attacks ($\sim$2%).
+LADE achieved 23% - 25% accuracy, REST 35% - 36%, and BiLD 23% - 25%, where success was defined as predicting a disease with similar symptoms (as per GPT-5 mini) compared to the ground truth, as temperature of the target model varies between 0.3 and 0.8 (see Table  3 in Appendix  F for detailed results).
+For comparison, we also perform a more “sophisticated” random guessing attack using OOD data: for each query in our test set, we randomly pick a disease from the OOD set as our guess, and check whether the guess has symptoms overlapping with the ground truth.
+This random guessing with OOD data has 6% accuracy (higher than 2% with our previous naive random guessing); this is still much lower than the 23% to 35% accuracy we achieve by training our classifier using the speculation traces observed with OOD data.
 These results show that although accuracy drops relative to the in-distribution setting, speculative decoding still leaks strong and learnable signals even under OOD conditions, without being much affected by temperature, underscoring the robustness of our side-channel attack.
@@
 In this section, we show how a malicious user can leak out private data used for predictions.
-We show this on REST He et al.
-(2024), which relies on retrieval from a datastore to generate speculative tokens.
+We show this on REST (He et al., 2024), which relies on retrieval from a datastore to generate speculative tokens.
 This datastore can be populated with proprietary information to tailor model output to specific domains, or data collected from user interactions, which may be leaked by our attack.
@@
 Leveraging Common Words.
-The LLM is prompted to generate paragraphs containing the most frequently occurring words, drawn from a dataset of the 10,000 most common English words Brants & Franz (2021).
+The LLM is prompted to generate paragraphs containing the most frequently occurring words, drawn from a dataset of the 10,000 most common English words (Brants & Franz, 2021).
 This exploits the high likelihood that phrases in the datastore contain these words.
@@
 […]
-No Padding Fixed Padding Random Padding (bytes) Packet Size Overhead
-(1024 bytes) D $\rightarrow$ 6 12 24 48 D $\rightarrow$ 6 12 24 48
-Experiment 1
+B.
+Random Padding.
+Packets can also be padded with a random number of bytes ($\epsilon$) selected from a uniform distribution of $\epsilon\sim Unif(0,D)$, where D is an upper bound of padding and $\epsilon$ is an integer.
 […]
-[…]
-[…]
-Experiment 2
-[…]
-[…]
-[…]
-Experiment 3
-[…]
-[…]
-[…]
-Table 4: Attack accuracy under constant size padding (packets padded to 1024 bytes) and variable-length padding (random padding) mitigations.
+Figure 6: Attack accuracy under variable-length padding (random padding) added to payloads within packets as a mitigation.
 Variable-length padding follows a uniform distribution, $\epsilon\sim Unif(0,D)$, with $D$ from 6 to 48.
 (Attack: 20 traces per prompt; model temperature: 0.8).
-[…]
 2.
@@
 This limits observability of fine-grain speculation patterns from packet sizes, reducing the success of fingerprinting attacks.
-Table 5 shows the impact of aggregating output tokens over 3 to 20 iterations.
-The accuracy of the exact knowledge attacks (Experiment 1 and 2) for LADE, REST, and BiLD drops to a minimum of 9.2%, 52%, and 11.2% respectively after aggregating tokens over 20 iterations down from 65%, 100%, and 49.2% respectively without aggregation; accuracy of the approximate knowledge attacks (Experiment 3) decreases to a minimum of 5.2%, 17.6%, and 4% respectively, down from 14%, 40%, and 20%.
+However, token aggregation linearly increases the inter-arrival time for packets, which linearly increases the latency observed by a user between text renderings that may hinder user experience.
+Figure 7 shows the accuracy for all our fingerprinting attacks after aggregating output tokens over 3 to 20 iterations.
+The accuracy of the exact knowledge attacks (Experiment 1 and 2) for LADE, REST, and BiLD drop to a minimum of 9.2%, 52%, and 11.2% respectively after aggregating tokens over 20 iterations, down from 65%, 100%, and 49.2% respectively without aggregation; accuracy of the approximate knowledge attacks (Experiment 3) decreases to a minimum of 5.2%, 17.6%, and 4% respectively, down from 14%, 40%, and 20% respectively.
 Thus, token aggregation is an effective mitigation.
@@
 At the extreme, aggregating all tokens in a single packet resembles a non-streaming LLM.
-Experiment 1 Experiment 2 Experiment 3
-AGG $\rightarrow$ 1 3 5 10 20 1 3 5 10 20 1 3 5 10 20
-LADE 65.6% 53.2% 40% 24.4% 10.8% 64.8% 45.2% 28.8% 16.4% 9.2% 13.6% 13.2% 14.4% 9.2% 5.2%
-REST 100% 96.8% 89.2% 80.4% 52% 100% 97.6% 92.8% 75.6% 52% 40% 33.2% 28.8% 20.4% 17.6%
-BiLD 49.2% 34% 30% 19.2% 11.2% 75.2% 57.6% 41.6% 30% 16.8% 20.4% 15.6% 10% 10% 4%
-Table 5: Token Aggregation Mitigation. Attack accuracy (20 traces/prompt, temperature of 0.8) as aggregation granularity (AGG), i.e., iterations tokens are aggregated over, varies from 3 to 20.
+Figure 7: Token Aggregation Mitigation.
+Attack accuracy (30 traces/prompt, temperature of 0.8) decreases as aggregation granularity (AGG) increases, i.e., iterations tokens are aggregated over, varies from 1 (no aggregation) to 20.
+3.
+Packet Splitting.
+Another way of limiting attacker observability of token counts, is to split the multiple tokens generated in an iteration into individual packets that are spread out in time throughout the decode iteration time period.
+To ensure that the attacker cannot observe even an approximate count of tokens-per-iteration by tracking the number of packets sent during one decode iteration, the defense needs to further also limit the number of packets per iteration and carry excess tokens to later iterations.
+This can introduce an overall latency impact, varying based on the framework, model and speculation technique.
+Future works can develop novel speculation techniques with such traffic shaping built in, to get both security and performance.
 6.2 Mitigating Datastore Leaks
@@
 This attack can be generalized to any value of parameter G in LADE.
-Appendix D Impact of Mitigation for Token-Length Side Channel
-In this experiment, we want to measure the impact on our attacks of mitigation against prior attacks, such as token-length side channels Weiss et al.
-(2024) that leaks the value of a token based on its associated character length.
+Appendix D Results for Random Padding Mitigation
+[…]
+No Padding Fixed Padding Random Padding (bytes) Payload Size Overhead
+(1024 bytes) D $\rightarrow$ 6 12 24 48 D $\rightarrow$ 6 12 24 48
+Experiment 1 - Exact Knowledge
+[…]
+[…]
+[…]
+Experiment 2 - Exact Knowledge (Constrained)
+[…]
+[…]
+[…]
+Experiment 3 - Approx Knowledge
+[…]
+[…]
+[…]
+Table 2: Attack accuracy under constant size padding (payload within packets padded to 1024 bytes) and variable-length padding (random padding) mitigations.
+Variable-length padding follows a uniform distribution, $\epsilon\sim Unif(0,D)$, with $D$ from 6 to 48.
+(Attack: 20 traces per prompt; model temperature: 0.8).
+Appendix E Impact of Mitigation for Token-Length Side Channel
+In this experiment, we want to measure the impact on our attacks of mitigation against prior attacks, such as token-length side channels (Weiss et al., 2024) that leaks the value of a token based on its associated character length.
 We evaluate the impact of a mitigation, where every token has been padded to the same size to mask its character length.
 We measure the attack success rate for our attacks with this mitigation.
-Note that even with this mitigation, there is variation in the packet sizes, due to simply the number of tokens contained in each packet, which varies due to speculation.
+Note that even with this mitigation, there is variation in the packet sizes, due to the number of tokens contained in each packet, which varies due to speculation.
 Thus, measuring the packet sizes in this scenario is the same as measuring the token counts.
+Figure 10 shows the accuracy for our attack despite this mitigation as temperature varies.
+At smaller temperatures, the accuracy of our attacks is still significant (with temperature 0.3, in both exact knowledge attacks (experiment 1 and 2), for LADE, the accuracy is about 40-50%, and for BiLD, it is about 80%).
+At high temperatures, our attack accuracy is reduced: with temperature 1.0, in both the exact knowledge scenarios (experiment 1 and 2), for LADE, the accuracy is reduced to around 20%, and for BiLD, it is reduced to 30%.
+Figure 10: Accuracy of the query fingerprinting attack with each token padded to the maximum size of tokens that can be generated per iteration (1024 bytes)
 […]
-Temperature
-0.3 0.6 0.8 1.0
-Experiment 1
-LADE 51.6% 38% 27.6% 22.8%
-REST 24% 17.2% 10.8% 8.4%
-BiLD 67.6% 43.2% 30.4% 22.8%
-Experiment 2
-LADE 41.2% 24.8% 22.4% 11.2%
-REST 25.6% 14% 14.4% 10%
-BiLD 82.4% 62.4% 46.8% 33.2%
-Experiment 3
-LADE 8% 6.8% 6.8% 3.2%
-REST 6.4% 4.8% 8% 5.2%
-BiLD 8.4% 10% 10.8% 6.8%
-Table 6: Accuracy of the Query Fingerprinting Attack with each token padded to the maximum token size (1024 bytes)
-Appendix E Out of Distribution Training
+Appendix F Out of Distribution Training
+F.1 Attack Accuracy
+Table 3 shows the attack accuracy using out of distribution dataset with 50 queries about diseases as training data, and using our original evaluation data set from Experiment 1 as the test set.
+Our attack achieves an accuracy of 23% to 36% of guessing overlapping symptoms by training wiht the OOD set, achieving significantly higher accuracy than random guessing using OOD set, which achieves only a accuracy of 6%.
 Temperature $\rightarrow$ 0.3 0.6 0.8
@@
 Table 3: Accuracy of the Query Fingerprinting Attack with out of distribution training, as the temperature of target model varies from 0.3 to 0.8.
+F.2 Degree of Overlap Between OOD Training Set and Evaluation Set
+Since we use a 50 class OOD data set (generated using ChatGPT) as training set, and a 50 class data set for the evaluation set (used in Experiment 1), there is no direct 1-1 correspondence between training and evaluation sets.
+However, artificially, if the symptoms of each disease in the OOD training set has a high overlap with all the diseases of the test set, then our attack accuracy can be artificially high.
+To understand whether this is the case, we study the the overlap in symptoms between diseases in the OOD training set and that in the evaluation set (Experiment 1 - Exact Knowledge).
+Figure 13 shows a 2D matrix illustrating symptom overlap between diseases in the OOD training set (X-axis) and diseases in the evaluation set from Experiment 1 (Y-axis).
+Both sets have 50 diseases.
+A cell is colored white if the corresponding pair of diseases shares overlapping symptoms.
+To determine overlap, we query an LLM (e.g., ChatGPT) with the two diseases and check whether they have common symptoms.
+In the OOD training set, a small number of diseases (3-4) share symptoms with multiple (5-6) diseases in the evaluation set, and vice versa.
+However, most training-set diseases overlap with only one or two evaluation-set diseases.
+Additionally, some evaluation-set diseases have no symptom overlap with any training-set disease; in such cases, an attacker cannot infer the patient’s symptoms.
+This is expected, as the OOD training set is independently constructed from the top 50 diseases most commonly queried on ChatGPT, while the evaluation set is derived from real-world user queries to medical chatbots.
+This diversity in overlap patterns reflects a realistic deployment setting. Consequently, the observed attack accuracy of 23%–36% is realistic.
+In comparison, a random guessing baseline attack, that assigns a random OOD training-set disease to each evaluation set disease achieves only 6% accuracy, aligning with our observation that most diseases in OOD training set do not overlap in symptoms with many test set diseases.
+This highlights the potency of our attack using speculative decoding when realistic OOD data is used for training.
+Figure 11: Overlap in symptoms between diseases in the OOD Training set (X-axis) and in the Evaluation set used in Experiment 1 - Exact Knowledge (Y-axis).
+A white entry indicates that two diseases have similar symptoms (as per GPT-4o), while a black entry indicates no overlap.
+Overall, there is minimal overlap in symptoms across any random pair of diseases from OOD training dataset and evaluation dataset, representative of realistic OOD datasets.
+Appendix G Comparing Classification Methods (Random Forest, GMM, CNN)
+In this section, we evaluate the accuracy of the query fingerprinting attack (Experiment 1 - Exact Knowledge) using different classifiers - Random Forest, GMM and CNN.
+We study this with for the three speculative decoding techniques LADE, REST and BiLD for a range of temperatures (0.3, 0.6, 0.8, 1), all with traces per query (TPQ) of 30.
+For the GMM classifier, we trained the scikit-learn’s GMM with 1 component per class, and make predictions via Bayes’ Rule.
+For the CNN classifier, we use three 1D convolution layers with max pooling followed by average pooling and a fully connected layer.
+For the Random Forest classifier, our setup is similar to Section 4.4.
+Figure 12: Accuracy of query fingerprinting attack (Experiment 1 - Exact Knowledge) using Random Forest, GMM, and CNN classifiers, for temperatures 0.3, 0.6, 0.8, 1.0 (using 30 traces per query).
+As shown in Figure 12, the attack accuracy for the GMM classifier ranges from 13-54% for LADE, 79-96% for REST, and 14-51% for BiLD.
+However the GMM performs worse than the Random Forest classifier which achieves an accuracy of 61-92% for LADE, 99-100% for REST, and 34-74% for BiLD.
+While prior work Carlini & Nasr (2024) showed that GMM can achieve query fingerprinting with high accuracy (up to 100%), we observe that their attacks were restricted to 2 classes, contrary to our attacks that classify across 50 classes, in which scenario the GMM achieves lower accuracy and the Random Forest is able to outperform it.
+In comparison, the CNN performs worse than both the Random Forest and the GMM, achieving an accuracy of 10-22% for LADE, 52-80% for REST, and 5-12% for BiLD.
+This is because the training data set is relatively small, with only 250 to 1500 training data points.
+So we observe that the CNN overfits to a given trace, leading to lower accuracy.
+These results lead us to choose Random Forest classifier for all our attacks in this paper, given that it outperforms GMM and CNN.
+Appendix H Scaling Behavior
+In this section, we study how the accuracy of the query fingerprinting attack (with Experiment 1 - Exact Knowledge) with temperature of 0.8 changes as the dataset size (number of diseases) scales from 5 to 50.
+As shown in Table 4, for BiLD and REST, the accuracy remains relatively stable as the number of diseases increase.
+The attack accuracy for LADE decreases by about 7 percent when we double the dataset size from 25 diseases to 50.
+This shows that the attack is relatively robust to scaling of the dataset size.
+Number of Diseases in Dataset
+5 10 15 20 25 30 35 40 45 50
+LADE 92% 86% 89.3% 84% 76.8% 74.7% 74.3% 73.5% 74.2% 68%
+REST 100% 100% 100% 100% 100% 100% 100% 99.5% 99.6% 99.6%
+BiLD 48% 40% 60% 58% 45.6% 48% 49.1% 48.5% 48.8% 52%
+Table 4: Accuracy of query fingerprinting attack (Experiment 1 - Exact Knowledge) with temperature of 0.8, using 30 traces per query (TPQ) for training as the number of diseases in the dataset scales from 5 to 50.
+Next, we study how the training cost scales with the dataset size.
+Specifically, we measure the cost in terms of the number of traces per query required to reach 90% of the fingerprinting attack (Experiment 1 - Exact Knowledge, with temperature 0.8, and dataset of 50 diseases).
+We measure this cost as the dataset size scales from 5 to 50 diseases.
+As shown in Table 5, we observe that the number of traces (cost) remains relatively stable (for REST) or increases moderately (LADE and BiLD).
+For REST, this is because the retrieval based speculation provides an extremely stable signal, which allows for high accuracy in the attack with a limited number of traces.
+In comparison, both LADE and BiLD have more variations, requiring more traces per query as the number of diseases increase.
+Number of Diseases in Dataset
+5 10 15 20 25 30 35 40 45 50
+LADE(61.2%) 3 3 3 4 5 7 7 10 9 9
+REST(89.6%) 4 2 3 4 4 5 5 3 4 4
+BiLD(46.8%) 22 30 15 16 28 29 28 27 19 19
+Table 5: Number of queries needed to achieve 90% of the fingerprinting attack accuracy (exact knowledge - experiment 1) with temperature of 0.8, as the number of diseases in the dataset scales from 5 to 50.
+Appendix I Ablation: Query Fingerprinting Attack on Insurance Chatbot
+To show that our attack is not restricted to the medical domain, and also works on other domains, we demonstrate our attack on an insurance chatbot.
+Figure 13: Attack accuracy of query fingerprinting with 50 insurance-related queries as temperatures vary (0.3, 0.6, 0.8, 1.0), using 30 Traces Per Query (TPQ) for training.
+We perform the query fingerprinting attack over a set of 50 queries consisting of insurance-related questions, generated by asking ChatGPT for commonly asked insurance-related questions by users.
+These include questions on health, car, and home insurance.
+Learning whether a user is asking such questions can allow an attacker to know private information such as whether a user owns a car or a home or has health concerns.
+We perform the Exact Knowledge, Exact (Constrained) and Approx Knowledge experiments , as designed for the medical chatbot in 4.3.
+Our REST and BiLD, our attack on an insurance chatbot achieves 97.2–100% and 37.6–91.6% query-fingerprinting accuracy for temperatures 0.3–1.0, slightly outperforming the results we observed on the medical chatbot.
+LADE with 17.2–59.2% attack accuracy performs worse than in the medical case, however it still outperforms random chance (2%).
+This shows that the attack does not rely on any medical specific terminologies and can generalize to more diverse domains.